### v1

# seq-JEPA: Autoregressive Predictive Learning of Invariant-Equivariant World Models

Abstract
Current self-supervised algorithms mostly rely on transformations such as data augmentation and masking to learn visual representations. This is achieved by inducing invariance or equivariance with respect to these transformations after encoding two views of an image. This dominant two-view paradigm can limit the flexibility of learned representations for downstream adaptation by creating performance trade-offs between invariance-related tasks such as image classification and more fine-grained equivariance-related tasks. In this work, we introduce seq-JEPA, a world modeling paradigm based on joint-embedding predictive architecture that leverages architectural inductive biases to resolve this trade-off. Without requiring an additional equivariance predictor or loss term, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to the specified transformations and another invariant to them and suited for tasks such as classification. To do so, our model processes a short sequence of different views (observations) of an input image. Each encoded view is concatenated with embeddings corresponding to the relative transformation (action) producing the next observation in the sequence. A transformer encoder outputs an aggregate representation of this sequence, which is subsequently conditioned on the action leading to the next observation to predict its representation. Empirically, seq-JEPA achieves strong performance on equivariant benchmarks and image classification without sacrificing one for the other. Additionally, our framework excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
Self-supervised learning, world models, computer vision
Figure 1: Top-1 linear classification (invariance) versus rotation prediction (equivariance) performance on 3DIEBench. Our world model, seqJEPA, learns representations for both invariance- and equivariance-related tasks at the same time. Subscripts indicate training and inference sequence lengths used in seq-JEPA.
1 Introduction
Self-supervised learning (SSL) in latent space has made significant progress in visual representation learning in recent years, closing the gap with supervised learning in many downstream tasks (van den Oord et al., 2019; Misra & van der Maaten, 2020; He et al., 2020; Chen et al., 2020; Grill et al., 2020; Chen & He, 2021; Caron et al., 2020, 2021; Zbontar et al., 2021; Bardes et al., 2022; Baevski et al., 2022; Assran et al., 2023). Most SSL methods learn their representations by comparing two transformed views of a sample image and inducing invariance with respect to these transformations (Chen et al., 2020; He et al., 2020; Dwibedi et al., 2021; HaoChen et al., 2021; Yeh et al., 2022; Caron et al., 2020, 2021; Ermolov et al., 2021; Assran et al., 2022; Zbontar et al., 2021; Bardes et al., 2022). Another group of methods employ techniques to preserve some information about the transformations and learn equivariant representations (Lee et al., 2021; Xiao et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024).
Equivariance is a required representational property for many fine-grained downstream tasks. For example, given representations that are invariant to color, it is not possible to distinguish between certain species of flowers or birds (Lee et al., 2021; Xiao et al., 2021). Additionally, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024). Although some of the equivariant SSL methods have empirically shown a slight performance improvement in downstream tasks traditionally associated with invariance (e.g., image classification) (Devillers & Lefort, 2022; Park et al., 2022; Gupta et al., 2023), recent studies have revealed a persistent trade-off between invariance and equivariance in self-supervised representation learning (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024). The existence of such a trade-off has been recently supported by theoretical arguments (Wang et al., 2024), highlighting the need for innovative architectural and objective design to mitigate it.
Humans and other animals rely on a sequence of actions and consequent observations (views) for developing appropriate visual representation during novel object learning  (Harman et al., 1999; Vuilleumier et al., 2002). For example, they recognize a 3-D object by changing their viewpoint and examining different sides of the object (Tarr et al., 1998). Inspired by this, we introduce seq-JEPA, a world modeling paradigm based on joint-embedding predictive architecture (LeCun, 2022; Dawid & LeCun, 2024) that leverages architectural inductive biases to enable such sequential processing in SSL and resolve the invariance-equivariance trade-off. Our model simultaneously learns two architecturally distinct representations: one equivariant to a specified set of transformations and another suited for invariance-related tasks, such as image classification.
Specifically, our framework (Figure 2) processes a short sequence of different views (observations) of an input image. Each encoded view is concatenated with embeddings corresponding to the relative transformation (action) that produces the next observation in the sequence. These concatenated representations are then fed to a transformer encoder, resembling working memory, that outputs an aggregate representation, which is subsequently conditioned on the action leading to the next observation and used to predict its representation.
Our results show that the encoded views in seq-JEPA become transformation/action-equivariant. Through ablation experiments, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network. In contrast, the aggregate representation of views, produced at the output of the sequence transformer, becomes largely transformation/action-invariant. This architectural disentanglement of invariant and equivariant representations in seq-JEPA is key to the competitive performance of our model compared to both invariant and equivariant SSL methods (Figure 1). We empirically demonstrate this advantage across multiple tasks and benchmarks. It is important to emphasize that, unlike many previous approaches in equivariant SSL (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), neither our objective function nor our architecture explicitly encourages the learning of invariant-equivariant representations. Instead, the learning of invariance and equivariance, along with their architectural separation, emerges as an implicit consequence of action-conditioned sequential predictive learning and the model architecture.
Furthermore, we explore other benefits of moving away from the dominant two-view paradigm in SSL and demonstrate that seq-JEPA is capable of solving tasks that inherently require aggregating a sequence of observations. In one scenario, inspired by embodied vision in primates, our model learns image representations without hand-crafted augmentations or masking, solely via predictive learning across simulated saccades (eye movements). Finally, seq-JEPA is also capable of path integration over a sequence of actions such as eye movements in the previously described scenario, or 3-D object rotations in 3DIEBench (Garrido et al., 2023). A summary of our contributions are as follows:
  • We propose seq-JEPA, a self-supervised world model that leverages a sequence of action-observation pairs to learn separate representations for equivariance- and invariance-demanding downstream tasks without requiring an additional equivariance predictor or loss term.
  • We empirically show the effectiveness of our world modeling paradigm in comparison with invariant and equivariant SSL methods on both invariance- and equivariance-related benchmarks without sacrificing one for the other.
  • We demonstrate that seq-JEPA is capable of solving tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
2 Method
2.1 Invariant and equivariant representations
Before describing the architecture and training procedure of seq-JEPA, we provide the nomenclature of invariant and equivariant representation learning in SSL (Dangovski et al., 2022; Devillers & Lefort, 2022). Consider $\mathcal{T}$ as a distribution of possible data transformations that is parametrized by a vector $t$. These transformations can be any type of hand-crafted data augmentation, such as geometric transformations or masking, that transform an original data point $x$ to different views. For example, we denote $x_{1}$ and $x_{2}$ as views generated from $x$ via two transformations $t_{1}$ and $t_{2}$ sampled from $\mathcal{T}$. Additionally, we denote $a$ as a transformation that transforms $x_{1}$ to $x_{2}$. Here, we distinguish between $t$ and $a$, and call $t$ a transformation and $a$ an action, i.e., a relative transformation that transforms $x_{1}$ to $x_{2}$. Furthermore, we denote $f$ as an encoder function that projects the data space into a latent space. Given a transformation $t$, the encoder and its representations are said to be equivariant to $t$ if
$\forall{t\in\mathcal{T}},~{}\exists{\,u_{t}}\;\mathrm{s.t.}\qquad f(t(x))=u_{t}(f(x)).$ (1)
We can also define equivariance given the action $a$ and two views $x_{1}$ and $x_{2}$ as
$\forall{a\in\mathcal{T}},~{}\exists{\,u_{a}}\;\mathrm{s.t.}\qquad f(x_{2})=u_{a}(f(x_{1})),$ (2)
where $u_{t}$ and $u_{a}$ are functions that produce displacements in latent space corresponding to their associated transformation/action. As a special case of equivariance, the encoder and its representations are said to be invariant to $t$ if $u_{t}$ and $u_{a}$ are identity functions, i.e., $f(t(x))=f(x)$ or $f(x_{2})=f(x_{1})$.
2.2 Architecture
The overall architecture of seq-JEPA is depicted in Figure 2. Denote $\{x_{i}\}_{i=1}^{M+1}$ as a sequence of views generated from a data sample $x$ using transformations $\{t_{i}\}_{i=1}^{M+1}$. The corresponding actions (relative transformations), whose embeddings are learned via a linear layer are denoted as $\{a_{i}\}_{i=1}^{M}$. A backbone encoder, $f_{\theta}$ encodes the first $M$ views and produces representations $\{z_{i}\}_{i=1}^{M}$. These representations, except for $z_{M}$, are concatenated with their corresponding action embeddings and fed to a sequence model, e.g., a transformer encoder, $g_{\phi}$ (no MLP projector is used after the encoder). This transformer module is tasked with aggregating the sequence of representations of action-observation pairs to produce an aggregate representation. We use a learnable token which we call [AGG] to learn the aggregate representation (this token resembles a [CLS] token in vision transformers (Dosovitskiy et al., 2020)). The output corresponding to [AGG] is therefore:
$z_{AGG}=g_{\phi}((z_{1},a_{1}),(z_{2},a_{2})...,(z_{M-1},a_{M-1}),z_{M})$ (3)
The representation $z_{AGG}$ is subsequently concatenated with the last action embedding $a_{M}$ which corresponds to the relative transformation between $x_{M}$ and $x_{M+1}$, and is fed to an MLP predictor module $h_{\psi}$ to predict the representation of $x_{M+1}$, i.e., $\hat{z}_{M+1}=h_{\psi}(z_{AGG},a_{M})$. The last view $x_{M+1}$, is itself encoded using a target encoder whose weights are an exponential moving average (EMA) of the encoder $f_{\theta}$ used to encode the sequence of observations. The representation $z_{M+1}$ is passed through an stop-gradient (sg) to prevent representational collapse. Our objective is then to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$, which we achieve by minimizing the following prediction loss,
$\mathcal{L}_{seq-JEPA}=1-\frac{{\hat{z}_{M+1}}}{\left\lVert{{\hat{z}_{M+1}}}\right\rVert_{2}}\cdot\frac{{\texttt{sg}(z_{M+1}})}{\left\lVert{{\texttt{sg}(z_{M+1}})}\right\rVert_{2}}.$ (4)
No additional equivariance predictor or loss function is used when training the framework.
Figure 2: Overview of the proposed architecture; seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn separate representations for equivariance- and invariance-related tasks.
2.3 Action and observation sets
In order to evaluate our model and measure its ability to generalize to different types of transformations, we consider three distinct sets of transformation and their corresponding action-observation pairs (Figure 3a). Descriptions of each transformation set are provided below, with additional details available in Section A.1.
3D Invariant Equivariant Benchmark (3DIEBench). We use the 3DIEBench dataset (Garrido et al., 2023) as our first evaluation benchmark. This dataset is specifically designed for measuring equivariance and provides 3D object renderings with different object rotations and floor hue and lighting angles and their corresponding parameters. We specifically study equivariance to the 3D rotation group ($SO(3)$) and further to floor and light hue in this benchmark.
Hand-crafted data augmentations. In this setting, observations are transformed image views generated by hand-crafted augmentations that are typically used in the two-view paradigm of SSL. The actions are relative augmentation parameters between two views. We measure equivariance to crop, color jitter, and blur transformations. As benchmarks for this setting, we use CIFAR100 and Tiny ImageNet datasets. To generate augmented views, we follow EquiMod’s augmentation strategy (Devillers & Lefort, 2022). For details on augmentations and the corresponding parameterization, see Section A.1.
Predictive learning across saccades. Going beyond transformations such as augmentations and rotation, we explore other ways of leveraging sequential observations to learn visual representations. Specifically, we designed a procedure to learn image representations without using any hand-crafted augmentations with a similar flavor to I-JEPA (Assran et al., 2023) but without dependence on masking. Our approach is enabled by seq-JEPA’s architecture and its ability to produce an aggregate representations from a sequence of observations. In this setting, given a full-size high-resolution image, we extract a number of smaller patches and use them as a sequence of partial observations. In our experiments with the STL-10 dataset, we use 32 by 32 patches, i.e. one-ninth of the original image.
The action in this setup corresponds to the relative position between the center of two subsequent patches, i.e. a simulated act of rapid eye movements (also known as saccades) from one point of the image to another. Although it is possible to use random positions as patch centers (fixations) for the sequence, we show that adopting two techniques inspired by embodied vision in primates improves the quality of the aggregate representation and its downstream classification performance (Figure 3b). The first technique utilizes image saliencies (Itti et al., 1998; Li, 2002; Zhaoping, 2014) to sample fixations. To generate saliency maps, we use a pre-trained saliency prediction model, DeepGaze IIE (Linardos et al., 2021). Given a dataset, we pre-extract and store the saliencies without any additional overhead during training. These saliency maps can be interpreted as how likely humans are to center their gaze (fixate) on a given pixel of the image. Therefore, sampling fixations from the probability distribution corresponding to saliency maps provides more informative patches, yet similar to uniform sampling, is prone to a high degree of redundancy and overlap between patches. To remedy this, i.e. minimize overlap among patches and avoid future fixation that revisit previously attended areas, we use another technique inspired by biological vision known as inhibition of return (IoR) (Posner et al., 1985). To implement IoR, after sampling a fixation from saliency maps, we zero out the sampling probabilities of an area with a diameter equal to the patch width surrounding the previous fixations before sampling a subsequent fixation. In this scenario, we are interested to see whether encoder representations become sensitive to fixation positions (spatial awareness), and therefore equivariance is measured with respect to fixation coordinates.
Figure 3: a. Types of transformations and observations used for training; seq-JEPA works with a variety of action-observation pairs such as 3D object rotations, hand-crafted augmentations, and simulated saccades. b. For predictive learning across saccades, image saliencies and inhibition of return help create more informative and less redundant patch sequences.
3 Related Work
Joint-embedding predictive architectures. In JEPAs (LeCun, 2022; Dawid & LeCun, 2024), a predictor asymmetry is introduced in one of the two branches that traditionally encode transformed views in SSL. The representation learning task is then framed as a world modeling objective (Ha & Schmidhuber, 2018; Hafner et al., 2019), i.e., the predictor is tasked with predicting the consequence of a data transformation (action) applied to the first view that results in the second view. The predictor is conditioned on transformation parameters such as positional masks to predict the latent representation of target blocks corresponding to the masks as in I-JEPA (Assran et al., 2023), or additionally on parameters corresponding to hand-crafted image augmentations as in Image World Model (IWM) (Garrido et al., 2024). Recent works have also explored JEPAs for physical understanding (Garrido et al., 2025) and reward-free offline planning (Sobal et al., 2025).
Non-generative world models. Non-generative world models predict the consequence of an action or transformation in latent space. A line of work considers such world models in contrastive SSL by predicting the representation of a transformed view from another via a contrastive objective and conditioned on the past or context (van den Oord et al., 2019; Gupta et al., 2024). The family of JEPAs defer the task of world modeling to a predictor module that is conditioned on the action (Assran et al., 2023; Garrido et al., 2024; Sobal et al., 2025). Non-generative world models have also garnered attention in model-based reinforcement learning (RL) and planning to improve sample efficiency (Schwarzer et al., 2021), to learn environment dynamics (Sobal et al., 2025), or to create an intrinsic novelty-based reward to improve exploration in self-predictive world models (Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023). A prominent example of such world models is BYOL-Explore (Guo et al., 2022), in which a recurrent cell follows the encoder and is conditioned on the current action in a planning task, followed by a predictor to capture the environment’s transition dynamics.
Equivariant representation learning. Many equivariant SSL methods extend invariant methods by introducing an additional equivariance objective. This objective can be direct prediction of parameters corresponding to augmented views (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024), or additional transformations such as rotation (Dangovski et al., 2022). EquiMod (Devillers & Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space in addition to their invariant objective. Xiao et al. (2021) use contrastive learning to train multiple projection heads for equivariance to each augmentation, where the views corresponding to the desired augmentation are considered as negative samples. In SEN (Park et al., 2022), similar to EquiMod, a predictor is used to predict the effect of transformations in latent space, yet without any invariance objective. ContextSSL (Gupta et al., 2024) learns representations that are equivariant only to the recent transformations in the current context by conditioning representations on both the current action and the past via a memory module, and further using a transformation prediction objective to avoid collapse to invariance. Another line of work including Shakerinava et al. (2022), CARE (Gupta et al., 2023), and Yerxa et al. (2024) do not assume access to transformation parameters yet require two pairs of views to which the same transformation is applied, and an equivariance loss function is minimized to induce equivariance in representations. Generative SSL methods based on masked image modeling (He et al., 2022; Bao et al., 2021; Xie et al., 2022) can become equivariant to position due to the use of positional embeddings. Finally, action-conditioned JEPAs condition the predictor on positional mask tokens (Assran et al., 2023) and augmentation parameters (Garrido et al., 2024) to induce equivariance in encoder representations without any additional equivariance objective.
Relationship to existing literature. Our framework belongs to the family of JEPAs and is a non-generative world model. We show that via architectural inductive biases, JEPAs can be leveraged to resolve the invariance-equivariance trade-off in SSL and enable processing of sequential action-observation pairs that is required to solve specific perception tasks. Unlike the majority of equivariant SSL methods discussed above, seq-JEPA does not use an equivariance predictor or loss function, nor does it require two pairs of views to which the same transformation is applied. We further highlight the differences between our framework and a recent equivariant method, ContextSSL. Specifically, seq-JEPA can be distinguished from ContextSSL on two aspects. First, unlike seq-JEPA, ContextSSL operates within the two-view SSL paradigm but uses a decoder-only transformer instead of the usual MLP projector so that its output embeddings are conditioned on the past. Second, while ContextSSL’s main goal is adapting equivariance to different transformations by alternating transformations in training data, our goal is to learn separate invariant and equivariant representations with respect to one or more transformations specified, and solve downstream tasks that require aggregating a sequence of action-observation pairs.
4 Experimental Setup
4.1 Compared methods and baselines
We compare seq-JEPA’s performance with both invariant and equivariant SSL methods. For invariant methods, we compare with SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022). For equivariant methods, we compare with SEN (Park et al., 2022), EquiMod (Devillers & Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024). Additionally, we consider another equivariant baseline which we call Conditional BYOL, where BYOL’s predictor is conditioned on the action that transforms the view of the target branch to the view of the online branch, with the aim that transformation information is also encoded in representations. This architecture is equivalent to a two-view version of seq-JEPA with a sequence length of one during both training and inference, and with no sequence aggregator on top of representations. Furthermore, as a baseline for predictive learning across saccades, we consider a model which we call Conv-JEPA. This baseline has the same architecture as Conditional BYOL, and to train it, we consider the same sequence of patches that seq-JEPA uses (sampled from image saliencies), predict the representation of the last patch from each individual patch in the sequence, and sum their prediction losses. For architectural details of each baseline, see Section A.3.
4.2 Training protocol
For all methods, we use a ResNet-18 (He et al., 2016) (not pretrained) as the backbone encoder. For experiments that use CIFAR100 and low-resolution STL-10 patches in predictive learning across saccades, we use the CIFAR variant of ResNet-18. For action conditioning, similar to EquiMod (Devillers & Lefort, 2022), we use a learnable linear projection to learn action embeddings. In our default setting for all experiments with action conditioning, the action projector is 128-dimensional. For the sequence aggregator, we use a small transformer encoder (Vaswani et al., 2017) with three layers, four attention heads, and post-normalization. For the predictor, we use an MLP with a hidden layer of 1024 dimensions and ReLU activation. In baseline comparisons, in order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train two variants of baselines that typically employ an MLP projector: one maintaining the original MLP projector and another where the projector is replaced with our transformer encoder. We report the higher performance among the two models. For a fair comparison, all models are trained for a fixed batch size of 512 and the same number of epochs. For 3DIEBench experiments, all models are trained for 1000 epochs, while for other datasets, they are trained for 2000 epochs. For models that use a transformer as projector, including seq-JEPA, we use the AdamW optimizer with default $\beta_{1}$ and $\beta_{2}$, a weight decay of $0.001$, and a learning rate of $4\times 10^{-4}$ with a linear warmup for $20$ epochs starting from $10^{-5}$ followed by a cosine decay back to $10^{-5}$. For models that use an MLP projector, we use the Adam optimizer with a learning rate of $10^{-3}$, default $\beta_{1}$ and $\beta_{2}$, and no weight decay.
4.3 Evaluation metrics and protocol
To measure equivariance with respect to a specific transformation, we perform action decoding from encoder representations as in Garrido et al. (2023). Specifically, given two transformed views, we extract representations from the frozen encoder, concatenate them, and train a regression head on top to predict the relative transformation (action) between the two. We report the $R^{2}$ score of this regressor on the test set as the equivariance metric. Although not always a perfect measure of invariance (Wang et al., 2024), we report top-1 classification accuracy of a linear probe as an approximate metric for representational invariance. For all baselines, we report the accuracy based on encoder representations. For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 2). When reporting accuracy based on $z_{AGG}$, we report the number of observation views (context lengths) used during both training and inference. For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 2) for all seq-JEPA models in Section B.4. Details on training the evaluation probes are given in Section A.2.
5 Results
5.1 Quantitative evaluation on 3DIEBench
We use the 3DIEBench benchmark to quantitatively compare invariant and equivariant representations in seq-JEPA with baseline methods. This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object category classification. Table 1 provides a summary of our evaluation results on the 3DIEBench dataset where equivariant methods have been conditioned on rotation. In addition to the relative rotation between two views, in the last column, we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view. For seq-JEPA, we trained models with training sequence lengths between one to four (denoted by $M_{tr}$ in the table) and measure the linear classification performance with different inference lengths (denoted by $M_{val}$) on top of the aggregate representation. Among the invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance. Among equivariant baselines, while EquiMod and SEN offer even better classification accuracy compared to invariant methods, in terms of equivariance, they lag behind SIE and ContextSSL which provide the best rotation prediction performance due to their specialized equivariance predictors and loss functions. Even though both SIE and ContextSSL representations become highly sensitive to rotation, they do not fare well in linear probe classification. Seq-JEPA is able to achieve a rotation $R^{2}$ on a par with the best performing equivariant method while outperforming all methods in classification thanks to sequence aggregation during training and inference with a margin that increases with the inference sequence length. Finally, when action conditioning is ablated (last row in the table), seq-JEPA loses its equivariant properties, yet maintains a strong classification performance. For additional results on 3DIEBench including models with different training and inference sequence lengths and models conditioned on rotation and color, see Sections B.1 to B.3.
Table 1: Quantitative evaluation of learned representations on 3DIEBench for rotation prediction (equivariance) and linear probe classification (invariance). Equivariant models and seq-JEPA are conditioned on rotation. For seq-JEPA, the training sequence length is denoted by $M_{tr}$ and probing on top of $z_{AGG}$ is done with an inference length $M_{val}$.
Method Classification (top-1) Relative Rotation ($R^{2}$) Individual Rotation ($R^{2}$)
Invariant
BYOL 82.90 0.12 0.25
SimCLR 81.13 0.35 0.54
VICReg 80.48 0.20 0.36
Equivariant
SEN 83.43 0.35 0.57
EquiMod 84.29 0.32 0.55
SIE 77.49 0.58 0.62
Conditional BYOL 82.61 0.31 0.47
ContextSSL ($c=126$) 80.4 0.74 0.78
Invariant-Equivariant
seq-JEPA ($M_{tr}=1$, $M_{val}=1$) 84.08 0.65 0.69
seq-JEPA ($M_{tr}=1$, $M_{val}=3$) 85.31 0.65 0.69
seq-JEPA ($M_{tr}=3$, $M_{val}=3$) 86.14 0.71 0.74
seq-JEPA ($M_{tr}=3$, $M_{val}=5$) 87.41 0.71 0.74
seq-JEPA (w/o act cond, $M_{tr,val}=3$) 86.05 0.29 0.37
5.2 Qualitative evaluation on 3DIEBench
In order to qualitatively evaluate the representations learned by seq-JEPA and shed some light on its invariance-equivariance properties, we conduct two different visualizations. First, to see equivariance in representational space, we retrieve the three nearest representations of a query image from the validation set of 3DIEBench (Figure 4). All methods successfully retrieve the correct object category, while only seq-JEPA and SIE, which demonstrate high quantitative equivariance performance (Table 1), maintain rotation equivariance across all three nearest representations.
Figure 4: Retrieval of nearest representations; given a query image, we extract the three nearest encoder representations in the validation set of 3DIEBench. The retrieved views of models with the highest quantitative rotation equivariance performance maintain the rotation of the query image.
In the second visualization, we plot the 2D UMAP projections of seq-JEPA’s encoder and aggregate representations (Figure 5) for the model trained conditioned on rotation. The left panel of of Figure 5 shows encoder representations colored by class label, while the middle panel displays the same encoder representations colored by rotation angle. The smooth color gradation across the map within each class cluster in this middle panel suggests that the encoder captures rotation angle as a continuous factor, implying equivariance to rotation (e.g., the red class in the bottom-right corner of the right panel and the corresponding part in the middle panel). The right panel shows aggregate representations colored by class label. Comparing the class-colored plots (left and right panels), we observe that both encoder and aggregate representations contain class information. However, when we aggregate multiple views of a sample, some of the intra-class variability (resulting from transformations such as rotation) is eliminated, causing each class’s representational cluster to become more homogeneous. This aggregation procedure likely reduces variation due to rotation and makes the representations more invariant, resulting in decreased intra-class spread and increased inter-class distance. We create a similar UMAP visualization for seq-JEPA with ablated rotation conditioning in Section B.5 which highlights the role of action conditioning in achieving equivariance to rotation.
Figure 5: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on 3DIEBench and conditioned on rotation with $M_{tr}=3$ and $M_{val}=5$. Encoder representations for each view observation, color-coded by class category (left) and rotation angle (middle). The aggregate token representation for $M_{val}=5$ observations during inference, color-coded by class category (right).
5.3 Evaluation of models trained using hand-crafted augmentations
We next quantitatively assess invariance and equivariance for the second set of transformations, i.e. hand-crafted augmentations. Table 2 provides a summary of our evaluation on models trained on CIFAR100 and Tiny ImageNet. For each equivariant method and seq-JEPA, we trained four different variations. In three variations, the model is conditioned on only one of the augmentations among cropping, color jittering, and blurring, and in the last variation, on all three augmentations. Considering equivariance to augmentations, seq-JEPA outperforms invariant and equivariant baselines by a large margin. Notably, except for the model trained on CIFAR-100 and conditioned on blur, the best equivariance performance for a given augmentation is achieved when the model is specialized and conditioned only on that augmentation. Furthermore, ablating actions (last row in the table) causes seq-JEPA to achieve a lower equivariance performance across transformations compared to action-conditioned models. In terms of linear probe classification accuracy on top of aggregate representations, our framework is competitive with both invariant and equivariant families, yet does not outperform all of them. For additional results with different training and inference sequence lengths for seq-JEPA models trained with augmented views, see Section B.3.
Table 2: Quantitative evaluation of learned representations with hand-crafted augmentations on CIFAR100 and Tiny ImageNet; equivariance is measured by predicting relative transformation parameters associated with crop, color jitter, or blur augmentations. For seq-JEPA, the training sequence length is denoted by $M_{tr}$ and probing is done with an inference length ($M_{val}$) of five.
CIFAR100 Tiny ImageNet
Conditioning Method Classification (top-1) Crop ($R^{2}$) Jitter ($R^{2}$) Blur ($R^{2}$) Classification (top-1) Crop ($R^{2}$) Jitter ($R^{2}$) Blur ($R^{2}$)
Invariant
- SimCLR 61.72 0.56 0.18 0.04 34.29 0.30 0.08 0.13
BYOL 62.17 0.39 0.05 -0.01 35.70 0.17 0.01 0.16
VICReg 61.35 0.49 0.11 -0.06 35.29 0.31 0.05 0.19
Equivariant
Crop+Jitter+Blur SEN 61.94 0.65 0.51 0.85 36.01 0.24 0.48 0.87
EquiMod 61.80 0.59 0.49 0.74 36.75 0.38 0.46 0.86
SIE 58.81 0.34 0.26 0.53 31.37 0.26 0.56 0.88
Conditional BYOL 60.63 0.56 0.46 0.73 35.49 0.34 0.54 0.87
Crop SEN 61.56 0.66 0.15 0.10 35.95 0.24 0.11 0.49
EquiMod 61.77 0.62 0.15 0.01 36.83 0.26 0.12 0.30
SIE 57.55 0.69 0.11 0.25 32.38 0.34 0.05 0.14
Conditional BYOL 60.17 0.55 0.10 -0.01 37.07 0.22 0.01 0.24
Color jitter SEN 61.78 0.50 0.52 0.02 36.59 0.26 0.50 0.21
EquiMod 61.53 0.44 0.50 -0.02 37.18 0.20 0.52 0.29
SIE 59.29 0.48 0.59 0.06 34.37 0.39 0.62 0.35
Conditional BYOL 61.30 0.36 0.52 0.02 37.93 0.24 0.62 0.30
Blur SEN 61.47 0.43 0.15 0.84 34.62 0.16 0.08 0.79
EquiMod 62.72 0.41 0.15 0.74 36.12 0.24 0.10 0.91
SIE 57.66 0.40 0.07 0.71 31.00 0.26 0.05 0.85
Conditional BYOL 60.74 0.36 0.11 0.69 36.17 0.19 0.02 0.85
Invariant-Equivariant
Crop+Jitter+Blur seq-JEPA ($M_{tr}=1$) 52.90 0.77 0.52 0.23 37.10 0.64 0.49 0.89
Crop+Jitter+Blur seq-JEPA ($M_{tr}=2$) 60.17 0.78 0.64 0.88 35.56 0.69 0.42 0.93
Crop+Jitter+Blur seq-JEPA ($M_{tr}=3$) 58.33 0.79 0.63 0.92 34.85 0.67 0.64 0.96
Crop seq-JEPA ($M_{tr}=2$) 59.32 0.78 0.01 0.10 35.74 0.70 0.12 0.46
Color Jitter seq-JEPA ($M_{tr}=3$) 58.62 0.68 0.68 0.29 35.21 0.60 0.66 0.62
Blur seq-JEPA ($M_{tr}=3$) 56.82 0.71 0.15 0.74 35.79 0.58 0.22 0.97
- seq-JEPA ($M_{tr}=2$) 58.37 0.64 0.14 0.16 35.97 0.52 0.18 0.47
5.4 Predictive learning across saccades
Quantitative evaluation. In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations alongside actions to learn visual representations. Table 3 summarizes our evaluation results for predictive learning across saccades on STL-10. While seq-JEPA only uses low-resolution patches with no augmentation, it achieves a classification performance comparable to SimCLR, an invariant model that is trained using strongly augmented image views with full resolution. Ablating action conditioning causes the accuracy on top of the aggregate representations to drop significantly, indicating that positional awareness is essential to forming semantic representations across simulated eye movements. Furthermore, comparison with the Conv-JEPA baseline highlights the importance of aggregating a sequence of patch representations in contrast to simply accumulating prediction errors between pairs of patches. Additionally, the last two ablation results in the table show that saliency-driven sampling of patches alongside IoR play a key role in sampling informative and non-overlapping patches and achieving high-quality aggregate representations. Interestingly, while random uniform patch sampling negatively impacts classification accuracy, it results in the highest positional equivariance, as the model samples patches and corresponding saccade actions from across the entire image and not just the salient regions.
Table 3: Quantitative evaluation of predictive learning across saccades on STL-10. Equivariance is measured with respect to the position of a fixated patch on the image. Seq-JEPA models use only low-resolution patches with no augmentation during training and inference. Unless stated otherwise, training and inference lengths ($M_{tr}$ and $M_{val}$) are set to four for all seq-JEPA variants.
Conditioning Classification (top-1) Position ($R^{2}$)
Invariant
- SimCLR (augmentations) 85.23 -
Equivariant
position Conv-JEPA ($M_{tr,val}=4$) 80.04 0.80
Invariant-Equivariant
- seq-JEPA 70.45 0.38
position seq-JEPA 83.44 0.80
position seq-JEPA ($M_{val}=6$) 84.12 0.80
position seq-JEPA (no saliency) 79.85 0.88
position seq-JEPA (no IoR) 77.97 0.85
Qualitative evaluation. Figure 6 shows UMAP projections for seq-JEPA representations trained on STL-10 via predictive learning across saccades. As for rotation in Figure 5, equivariance to fixation position can be observed from the bottom panels that show UMAP projection of encoder representations color-coded by X and Y coordinates. Similar to rotation in 3DIEBench, we provide a similar UMAP visualization for seq-JEPA with ablated position conditioning in Section B.5 for predictive learning across saccades to highlight the role of action conditioning.
Figure 6: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on STL-10 via predictive learning across saccades and conditioned on fixation position with $M_{tr,val}=4$. Encoder representations for each view observation, color-coded by class category (top-left) and $X$ and $Y$ pixel coordinates (bottom panels). The aggregate token representation for $M_{val}=4$ observation patches during inference, color-coded by class category (top-right).
5.5 Path integration across actions
While an agent executes a sequence of actions in an environment, transitioning from an initial state to a final state, it should be capable of tracking its position by integrating its own actions. This is also a crucial cognitive ability that enables animals to estimate their current state in their habitat (McNaughton et al., 2006). Here, we evaluate whether pre-trained seq-JEPA is able to solve path integration. Given the sequence of observations $\{x_{i}\}_{i=1}^{M+1}$ generated from transformations $\{t_{i}\}_{i=1}^{M+1}$ and the corresponding action embeddings $\{a_{i}\}_{i=1}^{M}$, we define the task of path integration over the sequence of actions as predicting the relative action that would directly transform $x_{1}$ to $x_{M+1}$ (which we denote as $a_{1,M}$) given $z_{AGG}$ and $a_{M}$. In other words, given the aggregate representation of a sequence of action-observation pairs and the next action, we would like to predict the overall position change from the starting point ($x_{1}$) to the end point ($x_{M+1}$). We consider path integration for rotation angles in 3DIEBench and across eye movements with STL-10. For rotations, the task is integrating a series of object rotations from the first view to the last, i.e. angular path integration. For eye movements, the task is integrating the eye movements from the first saccade to the last, i.e. visual path integration. To measure path integration performance, we train a regression head on top of the concatenation of $z_{AGG}$ and $a_{M}$ to predict $a_{1,M}$ for a given inference sequence length. Figure 7 shows the results for both angular and visual path integration. The red curve corresponds to the performance of the original seq-JEPA. The blue curve corresponds to experiments in which the action embeddings are ablated (they are set to zero during inference for all views). The green curve corresponds to experiments in which the encoder (visual) representations are ablated (they are set to zero when being concatenated with actions before being fed to the transformer). As expected, path integration becomes more difficult as the number of observations increases (red curves). Ablating action conditioning (blue curves) results in failure of path integration. On the other hand, ablating the visual representations (green curves) results only in a small performance drop compared to the original model, indicating that action conditioning is the key factor enabling path integration.
Figure 7: a) Visual path integration across eye movements, b) Angular path integration across object rotations (results over three random seeds)
5.6 Role of action conditioning
To better understand the importance of action conditioning in learning highly equivariant representations, we perform ablation experiments for seq-JEPA trained on 3DIEBench by changing the dimensionality of action embeddings and ablating actions from one of the two or both components of seq-JEPA, i.e. the transformer encoder and the MLP predictor. These ablation results are provided in Table 4. An important observation is that both predictor and transformer action conditioning are important in achieving the best equivariant performance, yet predictor action conditioning plays a more important role. Furthermore, learnable action embeddings play a key role in developing equivariance to rotation, although this equivariance shows little sensitivity to the dimensionality of the action embedding. Additionally, classification performance is slightly boosted in the presence of action conditioning, yet its role in this case is not as important compared to positional embeddings in predictive learning across saccades.
Table 4: Ablation results on the effect of seq-JEPA’s action conditioning trained on 3DIEBench (equivariance to rotation); we investigate the effects of partial or complete action ablations and changing the dimensionality of action embeddings on linear classification and rotation prediction performance. All models have $M_{tr}=3$ and $M_{val}=5$ and results are reported over three randoms seeds.
Variant Classification (top-1) Rotation ($R^{2}$)
Act. conditioning
None 87.36 $\pm$ 0.7 0.29 $\pm$ 0.04
Random noise 58.16 $\pm$ 3.2 0.20 $\pm$ 0.07
Ablated predictor act cond. 87.17 $\pm$ 0.3 0.37 $\pm$ 0.06
Ablated transformer act cond. 86.33 $\pm$ 0.1 0.53 $\pm$ 0.05
Act. embedding dim
Non-learnable act emb 85.68 $\pm$ 0.2 0.34 $\pm$ 0.04
$adim=16$ 86.29$\pm$ 0.4 0.70 $\pm$ 0.01
$adim=64$ 87.11$\pm$ 0.2 0.70 $\pm$ 0.02
$adim=128$ (default) 87.41 $\pm$ 0.5 0.71 $\pm$ 0.02
$adim=256$ 87.26$\pm$ 0.6 0.72 $\pm$ 0.00
Figure 8: Effect of training and inference sequence length on seq-JEPA’s performance; left:: Equivariant performance ($R^{2}$) for different training sequence lengths (inference sequence length fixed); middle: Classification performance on top of aggregate representations versus training sequence length (inference sequence length fixed); right: Classification performance on top of aggregate representations for different inference sequence lengths (training sequence length fixed)
5.7 Role of training and inference sequence lengths
In order to study the effect of training and inference sequence lengths on our framework’s performance, we trained seq-JEPA with different training sequence lengths and measured both equivariance of encoder representations and linear probe accuracy using different inference sequence lengths to form the aggregate representation (Figure 8). A general positive trend can be observed in equivariance performance for different transformations when increasing the training sequence length (Figure 8, left panel), which can be the impact of seeing more action-conditioned representations during training. For linear classification (invariance-related performance; Figure 8, middle panel), the models trained on 3DIEBench and via predictive learning across saccades benefit from a longer training sequence length. However, in the augmentation setting with CIFAR100, a small drop in accuracy is observed for higher training sequence lengths. The observed differences across evaluation tasks likely stem from specific task requirements. We hypothesize that leveraging and aggregating a sequence of action-observation pairs, i.e. seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observation. In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased sequence length. Similarly, in the case of predictive learning across saccades, each eye movement and its subsequent glance provides additional information that can be leveraged for learning a richer aggregate representation. Such an advantage may not be present when using artificially-augmented images which can also explain the classification performance gap between seq-JEPA and the best-performing baselines in this setting (Table 2). Finally, during inference (Figure 8, right panel), all transformation settings benefit from more views and a richer context to improve the quality of aggregate representation.
6 Limitations and Future Perspectives
In this paper, we proposed seq-JEPA, a world modeling paradigm for learning architecturally segregated transformation-invariant and -equivariant representations. Although we have shown the feasibility of our world model for multiple transformation settings in image modality, we have not yet explored other modalities such as video, text, or multi-modal data. Given the architecture of seq-JEPA, it is possible to utilize the sequence aggregator to fuse representations from multiple modalities. Another promising direction is to build efficient world models for embodied agents with egocentric vision via scaling our approach to larger image and video datasets and take advantage of a sequence of foveated observations based on salience in a setup similar to predictive learning across saccades. Our transfer learning results on ImageNet-1k in Section B.6 hints at the possibility of building such a world model. Finally, similar to other equivariant SSL techniques, seq-JEPA requires knowledge of transformation groups. Developing a group-agnostic equivariant method is a challenge that remains largely unexplored in SSL.
Acknowledgements
This project was supported by funding from NSERC (Discovery Grants RGPIN-2022-05033 to E.B.M., and RGPIN-2023-03875 to S.B.), Canada CIFAR AI Chairs Program and Google to E.B.M., Canada Excellence Research Chairs (CERC) Program, Mila - Quebec AI Institute, Institute for Data Valorization (IVADO), CHU Sainte-Justine Research Centre, Fonds de Recherche du Québec–Santé (FRQS), and a Canada Foundation for Innovation John R. Evans Leaders Fund grant to E.B.M. This research was also supported in part by Digital Research Alliance of Canada (DRAC) and Calcul Québec.
Appendix A Implementation Details
A.1 Data preparation
3DIEBench. The original $256\times 256$ images are resized to a $128\times 128$ resolution for all experiments. Normalization is done using the means and standard deviations provided by Garrido et al. (2023), i.e., $\mu=[0.5016,0.5037,0.5060]$ and $\sigma=[0.1030,0.0999,0.0969]$ for the three RGB channels, respectively.
CIFAR100. We use $32\times 32$ images with normalization parameters typically used in the literature, i.e., $\mu=[0.4914,0.4822,0.4465]$ and $\sigma=[0.247,0.243,0.261]$. For data augmentation, we follow EquiMod’s augmentation strategy.
Tiny ImageNet. The training set consists of 100000 ImageNet-1k images from 200 classes (500 for each class) downsized to $64\times 64$. The validation set has 50 images per class. We use normalization parameters $\mu=[0.4914,0.4822,0.4465]$ and $\sigma=[0.247,0.243,0.261]$, for the three RGB channels, respectively. For data augmentation, we use the same augmentation parameters as CIFAR100 with the kernel size of Gaussian blur adapted to the $64\times 64$ images.
STL10. In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE (Linardos et al., 2021), resize the output saliencies back to $96\times 96$, and store them alongside original images. We use normalization parameters $\mu=[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively. After sampling fixations from saliencies, the patches are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$).
Transformation parameters. For 3DIEBench, we use the rotation and color parameters provided with images for action conditioning as done in (Garrido et al., 2023). For the augmentation setting, we use the parameters corresponding to each of the three augmentations and form the action as the relative augmentation vector between two images. For crop, we use four variables, i.e., vertical and horizontal coordinate, and height and width. For color jitter, we use four variables: brightness, contrast, saturation, and hue. For blur, we use one variable: the standard deviation of the blurring kernel. In predictive learning across saccades, the action is a 2-d vector, i.e., the normalized relative $(x,y)$ coordinate between two patches.
A.2 Training and evaluation details
Protocols for training evaluation heads.
For linear probing, we follow a common SSL protocol and train a linear classifier on top of frozen representations with a batch size of 256 for 300 epochs using the Adam optimizer with default hyperparameters. For action prediction, we follow a similar protocol as SIE (Garrido et al., 2023). Specifically, for rotation, color jitter, and crop, we train an MLP regressor with a hidden dimension of 1024 and ReLU activation for 300 epochs. For color (in 3DIEBench), blur (in the augmentation setting), and position (in predictive learning across saccades), we use a linear regressor and train it for 50 epochs. For path integration experiments, the same regressor architectures as the equivariance evaluation heads are used, i.e., the MLP for angular path integration and the linear regressor for saccade path integration. All regression heads are trained using the Adam optimizer with default hyperparameters.
Hardware. Each experiment was run on a single NVIDIA A100 GPU with 40GB of accelerator RAM. A single run of seq-JEPA on 3DIEBench with a training sequence length of three takes around 15 hours.
A.3 Architectural details
Below, we describe the architectural details and hyperparameters specific to each baseline.
BYOL.
We use a a projection head with 2048-2048-2048 intermediate dimensions. The predictor has a hidden dimension of 512-d with ReLU activation. We use the same EMA setup outlined in the original paper (Grill et al., 2020), i.e., the EMA parameter $\tau$ starts from $\tau_{\text{base}}=0.996$ and is increased following a cosine schedule.
SimCLR.
We use a temperature parameter of $\tau=0.5$ with a projection MLP with 2048-2048-2048 intermediate dimensions.
VICReg.
We use $\lambda_{\text{inv}}=\lambda_{V}=10$, $\lambda_{C}=1$, and a projection head with 2048-2048-2048 intermediate dimensions.
Conditional BYOL.
The architecture is the same as BYOL, except that the predictor also receives the normalized relative transformation parameters. We use a linear action projector of 128-d and the same EMA setup as BYOL.
SIE.
For both invariant and equivariant projection heads, we use intermediate dimensions of 1024-1024-1024. For the loss coefficients, we use $\lambda_{\text{inv}}=\lambda_{V}=10$, $\lambda_{\text{equi}}=4.5$, and $\lambda_{C}=1$. We use the hypernetwork architecture for all experiments.
SEN.
We use a temperature parameter of $\tau=0.5$ with a projection MLP with 2048-2048-2048 intermediate dimensions.
EquiMod.
We use the version based on SimCLR (both invariant and equivariant losses are contrastive with $\tau=0.1$ and have equal weights). The projection head has 1024-1024-128 intermediate dimensions. We use a linear action projector of 128-d.
ContextSSL.
We use the pre-trained weights provided by the authors (trained for 1000 epochs) and follow their evaluation protocol on 3DIEBench.
seq-JEPA.
For the sequence aggregator, we use a transformer encoder with three layers, four attention heads, and post-normalization. For the predictor, we use an MLP with a hidden layer of 1024-d and ReLU activation. The linear action projector in our default setting is 128-d. We use the same EMA setup as BYOL.
Appendix B Additional Experimental Results
B.1 Evaluation results on 3DIEBench for models conditioned on rotation and color
We further provide evaluation results for seq-JEPA and equivariant baselines on 3DIEBench conditioned on both rotation and color transformations in Table 5. The results show that all methods suffer a drop in classification performance and become highly sensitive to color. Similar performance degradations have been previously observed with 3DIEBench (Garrido et al., 2023; Gupta et al., 2024), yet without a convincing explanation. One possible explanation in line with Principle II in Wang et al. (2024) is that in 3DIEBench, floor and light hue are fairly decoupled from class information (low class relevance). Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
Table 5: Quantitative evaluation of learned representations on 3DIEBench for rotation and color prediction (equivariance) and linear probe classification (invariance). All models are conditioned on both rotation and color.
Method Classification (top-1) Rotation pred. ($R^{2}$) Color pred. ($R^{2}$)
SEN 82.17 0.29 0.96
EquiMod 82.58 0.27 0.95
Conditional BYOL 81.95 0.38 0.94
SIE 75.34 0.46 0.97
seq-JEPA ($M_{tr}=4$, $M_{val}=5$) 79.31 0.52 0.97
B.2 Measuring invariance of aggregate representations
In order to have a sense of invariance of the aggregate representation, we tried to predict the individual rotation parameters of each view for a seq-JEPA model with $M_{val}=3$ trained on 3DIEBench and conditioned on rotation. To avoid leaking the rotation parameters to $z_{AGG}$ during prediction, we zero out the action embeddings. The mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; which indicates that the aggregate representation has not encoded rotation in any of the views.
B.3 Complete evaluation results for linear probing on top of aggregate representations
In this section, we provide our complete evaluation results for linear probing on top of seq-JEPA’s aggregate representations for three transformation sets with different training and inference sequence lengths. Figure 9 shows the top-1 accuracy on 3DIEBench models conditioned on rotation. Figure 10 shows top-1 accuracy on STL-10 for models trained via predictive learning across saccades. Figures 11 and 12 show top-1 accuracy on CIFAR100 and Tiny ImageNet, respectively, with different types of action conditioning (crop, color jitter, blur, or all three). We can see the general trends of Figure 8 discussed in Section 5.7 also in these heatmaps.
Figure 9: seq-JEPA’s performance on 3DIEBench with rotation conditioning; the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
Figure 10: Seq-JEPA’s performance on STL-10 with predictive learning across saccades; the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
Figure 11: Seq-JEPA’s performance on CIFAR100 with different types of action conditioning (crop, color jitter, blur, or all three); the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
Figure 12: Seq-JEPA’s performance on Tiny ImageNet with different types of action conditioning (crop, color jitter, blur, or all three); the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
B.4 Comparison of evaluation results on encoder representations and aggregate representations
For completeness, we provide linear probe classification on encoder representations for different transformation settings in Table 6 and compare them with accuracy on aggregate representations for different inference evaluation lengths. The aggregate representation generally achieves a higher classification performance thanks to the architectural inductive bias in seq-JEPA.
Table 6: Comparison of seq-JEPA classification performance across datasets and conditioning. Top-1 classification accuracy is reported for $z_{res}$ and $z_{agg}$, with varying inference lengths $M_{eval}$.
Dataset Conditioning $M_{tr}$ $z_{res}$ $z_{agg}$
$M_{eval}=1$ $M_{eval}=3$ $M_{eval}=5$
3DIEBench None 3 80.91 81.61 86.05 87.36
3DIEBench Rotation 1 84.88 84.08 85.34 85.31
3DIEBench Rotation 3 82.49 81.72 86.14 87.41
3DIEBench Rotation + Color 4 74.88 71.14 75.97 79.31
CIFAR100 None 2 53.00 51.60 57.05 58.37
CIFAR100 Crop + Jitter + Blur 1 56.23 54.24 52.90 52.92
CIFAR100 Crop + Jitter + Blur 2 52.07 53.07 59.34 60.35
CIFAR100 Crop + Jitter + Blur 3 46.31 49.48 57.60 58.33
CIFAR100 Crop 2 52.62 52.06 58.07 59.32
CIFAR100 Color Jitter 3 54.92 50.21 57.20 58.62
CIFAR100 Blur 3 51.41 48.69 55.54 56.72
Tiny ImageNet None 2 32.84 30.48 35.03 35.97
Tiny ImageNet Crop + Jitter + Blur 1 33.03 32.34 36.14 37.07
Tiny ImageNet Crop + Jitter + Blur 2 27.20 30.57 34.99 35.56
Tiny ImageNet Crop + Jitter + Blur 3 24.74 28.84 33.78 34.68
Tiny ImageNet Crop 2 31.13 30.89 35.07 35.69
Tiny ImageNet Color Jitter 3 31.85 29.05 34.27 35.21
Tiny ImageNet Blur 3 27.20 28.83 34.82 35.79
STL-10 None 4 61.38 62.21 69.06 71.12
STL-10 Position 4 81.20 71.45 81.53 83.67
STL-10 Position (no saliency) 4 79.29 63.14 76.93 80.09
STL-10 Position (no IoR) 4 72.49 68.95 76.84 78.21
B.5 Visualization of UMAP projections for models without action conditioning
In Figure 13 we provide the UMAP projections for seq-JEPA representations trained on 3DIEBench but without action conditioning. Figure 14 shows UMAP projections for seq-JEPA representations trained on STL-10 via predictive learning across saccades without action conditioning. In these visualizations, the color gradients that implied high equivariance in UMAPs of action-conditioned models are either absent or less pronounced.
Figure 13: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on 3DIEBench without any action conditioning with $M_{tr}=3$ and $M_{val}=5$. Encoder representations for each view observation, color-coded by class category (left) and rotation angle (middle). The aggregate token representation for $M_{val}=5$ observations during inference, color-coded by class category (right).
Figure 14: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on STL-10 via predictive learning across saccades with no action conditining and with $M_{tr,val}=4$. Encoder representations for each view observation, color-coded by class category (left) and $X$ and $Y$ pixel coordinates (middle panels). The aggregate token representation for $M_{val}=4$ observation patches during inference, color-coded by class category (right).
B.6 Transfer learning results on ImageNet-1k
In this section, we consider transfer learning on ImageNet-1k using our frozen model trained on STL-10 via predictive learning across saccades. We follow the same evaluation protocol as the in-distribution setting and train a linear probe on ImageNet-1k training set and evaluate it on the validation set using the aggregate representation for different inference sequence lengths. We use DeepGaze-IIE to extract saliency maps of ImageNet-1k, resize them to $224\times 224$, and use them to extract the foveated patches. Figure 15 shows the top-1 accuracy on ImageNet-1k for two different patch sizes (32 and 84). These results hint at the possibility of scaling seq-JEPA to larger data regimes by striking a trade-off between the patch size and sequence length. For example, using full-resolution $224\times 224$ images, a forward pass through a ResNet-18 encoder takes approximately $1.83\times 10^{9}$ flops, while a forward pass of seq-JEPA with the default transformer encoder, $M_{val}=4$ and $84\times 84$ patches takes $1.25\times 10^{9}$ flops. Using the CIFAR variant of ResNet-18 with $32\times 32$ patches and $M_{val}=8$ requires $2.35\times 10^{9}$ flops.
Figure 15: Linear probe transfer learning accuracy on ImageNet-1k for two different patch sizes; the model is trained on STL-10 via predictive learning across saccades.

### v2

@@
 Abstract
-Current self-supervised algorithms mostly rely on transformations such as data augmentation and masking to learn visual representations.
-This is achieved by inducing invariance or equivariance with respect to these transformations after encoding two views of an image.
-This dominant two-view paradigm can limit the flexibility of learned representations for downstream adaptation by creating performance trade-offs between invariance-related tasks such as image classification and more fine-grained equivariance-related tasks.
-In this work, we introduce seq-JEPA, a world modeling paradigm based on joint-embedding predictive architecture that leverages architectural inductive biases to resolve this trade-off.
-Without requiring an additional equivariance predictor or loss term, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to the specified transformations and another invariant to them and suited for tasks such as classification.
-To do so, our model processes a short sequence of different views (observations) of an input image.
-Each encoded view is concatenated with embeddings corresponding to the relative transformation (action) producing the next observation in the sequence.
-A transformer encoder outputs an aggregate representation of this sequence, which is subsequently conditioned on the action leading to the next observation to predict its representation.
-Empirically, seq-JEPA achieves strong performance on equivariant benchmarks and image classification without sacrificing one for the other.
-Additionally, our framework excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
-Self-supervised learning, world models, computer vision
-Figure 1: Top-1 linear classification (invariance) versus rotation prediction (equivariance) performance on 3DIEBench.
-Our world model, seqJEPA, learns representations for both invariance- and equivariance-related tasks at the same time.
-Subscripts indicate training and inference sequence lengths used in seq-JEPA.
+Current self-supervised algorithms commonly rely on transformations such as data augmentation and masking to learn visual representations.
+This is achieved by enforcing invariance or equivariance with respect to these transformations after encoding two views of an image.
+This dominant two-view paradigm often limits the flexibility of learned representations for downstream adaptation by creating performance trade-offs between high-level invariance-demanding tasks such as image classification and more fine-grained equivariance-related tasks.
+In this work, we proposes seq-JEPA, a world modeling framework that introduces architectural inductive biases into joint-embedding predictive architectures to resolve this trade-off.
+Without relying on dual equivariance predictors or loss terms, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to specified transformations and another invariant to them.
+To do so, our model processes short sequences of different views (observations) of inputs.
+Each encoded view is concatenated with an embedding of the relative transformation (action) that produces the next observation in the sequence.
+These view-action pairs are passed through a transformer encoder that outputs an aggregate representation.
+A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation.
+Empirically, seq-JEPA demonstrates strong performance on both equivariant and invariant benchmarks without sacrificing one for the other.
+Furthermore, it excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
 1 Introduction
-Self-supervised learning (SSL) in latent space has made significant progress in visual representation learning in recent years, closing the gap with supervised learning in many downstream tasks (van den Oord et al., 2019; Misra & van der Maaten, 2020; He et al., 2020; Chen et al., 2020; Grill et al., 2020; Chen & He, 2021; Caron et al., 2020, 2021; Zbontar et al., 2021; Bardes et al., 2022; Baevski et al., 2022; Assran et al., 2023).
-Most SSL methods learn their representations by comparing two transformed views of a sample image and inducing invariance with respect to these transformations (Chen et al., 2020; He et al., 2020; Dwibedi et al., 2021; HaoChen et al., 2021; Yeh et al., 2022; Caron et al., 2020, 2021; Ermolov et al., 2021; Assran et al., 2022; Zbontar et al., 2021; Bardes et al., 2022).
-Another group of methods employ techniques to preserve some information about the transformations and learn equivariant representations (Lee et al., 2021; Xiao et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024).
-Equivariance is a required representational property for many fine-grained downstream tasks.
+Self-supervised learning (SSL) in latent space has made significant progress in visual representation learning, closing the gap with supervised methods across many tasks.
+Most SSL methods rely on comparing two transformed views of an image and enforcing invariance to the transformations (Misra and van der Maaten, 2020; Chen et al., 2020; He et al., 2020; Dwibedi et al., 2021; HaoChen et al., 2021; Yeh et al., 2022; Caron et al., 2020, 2021; Ermolov et al., 2021; Assran et al., 2022; Zbontar et al., 2021; Bardes et al., 2022).
+Another group of methods employ techniques to preserve transformation-specific information, thereby learning equivariant representations (Lee et al., 2021; Xiao et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023, 2024; Gupta et al., 2023, 2024; Yerxa et al., 2024).
+Equivariance is a crucial representational property for downstream tasks that require fine-grained distinctions.
 For example, given representations that are invariant to color, it is not possible to distinguish between certain species of flowers or birds (Lee et al., 2021; Xiao et al., 2021).
-Additionally, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024).
-Although some of the equivariant SSL methods have empirically shown a slight performance improvement in downstream tasks traditionally associated with invariance (e.g., image classification) (Devillers & Lefort, 2022; Park et al., 2022; Gupta et al., 2023), recent studies have revealed a persistent trade-off between invariance and equivariance in self-supervised representation learning (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024).
-The existence of such a trade-off has been recently supported by theoretical arguments (Wang et al., 2024), highlighting the need for innovative architectural and objective design to mitigate it.
-Humans and other animals rely on a sequence of actions and consequent observations (views) for developing appropriate visual representation during novel object learning  (Harman et al., 1999; Vuilleumier et al., 2002).
+Moreover, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024).
+While some equivariant SSL approaches have reported minor gains on tasks typically associated with invariance (e.g., classification) (Devillers and Lefort, 2022; Park et al., 2022; Gupta et al., 2023), a growing body of work highlights a fundamental trade-off between learning invariance and equivariance, i.e., models that capture equivariance-related style latents do not fare well in classification and vice versa (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024; Rusak et al., 2025).
+This trade-off has recently received theoretical support (Wang et al., 2024), underscoring the need for new architectural and objective designs that can reconcile these competing goals.
+In contrast to the two-view paradigm in SSL, humans and other animals rely on a sequence of actions and consequent observations (views) for developing appropriate visual representation during novel object learning  (Harman et al., 1999; Vuilleumier et al., 2002).
 For example, they recognize a 3-D object by changing their viewpoint and examining different sides of the object (Tarr et al., 1998).
-Inspired by this, we introduce seq-JEPA, a world modeling paradigm based on joint-embedding predictive architecture (LeCun, 2022; Dawid & LeCun, 2024) that leverages architectural inductive biases to enable such sequential processing in SSL and resolve the invariance-equivariance trade-off.
-Our model simultaneously learns two architecturally distinct representations: one equivariant to a specified set of transformations and another suited for invariance-related tasks, such as image classification.
-Specifically, our framework (Figure 2) processes a short sequence of different views (observations) of an input image.
-Each encoded view is concatenated with embeddings corresponding to the relative transformation (action) that produces the next observation in the sequence.
-These concatenated representations are then fed to a transformer encoder, resembling working memory, that outputs an aggregate representation, which is subsequently conditioned on the action leading to the next observation and used to predict its representation.
-Our results show that the encoded views in seq-JEPA become transformation/action-equivariant.
-Through ablation experiments, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network.
-In contrast, the aggregate representation of views, produced at the output of the sequence transformer, becomes largely transformation/action-invariant.
-This architectural disentanglement of invariant and equivariant representations in seq-JEPA is key to the competitive performance of our model compared to both invariant and equivariant SSL methods (Figure 1).
-We empirically demonstrate this advantage across multiple tasks and benchmarks.
-It is important to emphasize that, unlike many previous approaches in equivariant SSL (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), neither our objective function nor our architecture explicitly encourages the learning of invariant-equivariant representations.
-Instead, the learning of invariance and equivariance, along with their architectural separation, emerges as an implicit consequence of action-conditioned sequential predictive learning and the model architecture.
-Furthermore, we explore other benefits of moving away from the dominant two-view paradigm in SSL and demonstrate that seq-JEPA is capable of solving tasks that inherently require aggregating a sequence of observations.
-In one scenario, inspired by embodied vision in primates, our model learns image representations without hand-crafted augmentations or masking, solely via predictive learning across simulated saccades (eye movements).
-Finally, seq-JEPA is also capable of path integration over a sequence of actions such as eye movements in the previously described scenario, or 3-D object rotations in 3DIEBench (Garrido et al., 2023).
-A summary of our contributions are as follows:
-  • We propose seq-JEPA, a self-supervised world model that leverages a sequence of action-observation pairs to learn separate representations for equivariance- and invariance-demanding downstream tasks without requiring an additional equivariance predictor or loss term.
-  • We empirically show the effectiveness of our world modeling paradigm in comparison with invariant and equivariant SSL methods on both invariance- and equivariance-related benchmarks without sacrificing one for the other.
-  • We demonstrate that seq-JEPA is capable of solving tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
+Inspired by this, we introduce seq-JEPA, a self-supervised world modeling framework that combines joint-embedding predictive architectures (LeCun, 2022) with inductive biases for sequential processing.
+seq-JEPA simultaneously learns two architecturally distinct representations: one that is equivariant to a specified set of transformations, and another that is suited for invariance-demanding tasks, such as image classification.
+Specifically, our framework (Figure 1) processes a short sequence of transformed views (observations) of an image.
+Each view is encoded and concatenated with an embedding corresponding to the relative transformation (action) that produces the next observation in the sequence.
+These view-action pairs are passed through a transformer encoder, a form of learned working memory, that outputs an aggregate representation of them.
+A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation.
+Our results demonstrate that individual encoded views in seq-JEPA become transformation/action-equivariant.
+Through ablations, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network.
+In contrast, the aggregate representation of views, produced at the output of the transformer, becomes largely transformation/action-invariant.
+This emergent architectural disentanglement of invariant and equivariant representations is central to seq-JEPA’s competitive performance compared to invariant and equivariant SSL methods on both categories of tasks (Figure 3).
+Unlike most prior equivariant SSL methods (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), our model does not rely on explicitly crafted loss terms or objectives to achieve equivariance, nor is it instructed to learn the decomposition of two representation types.
+Instead, the dual representation structure arises naturally from the model architecture and action-conditioned predictive learning.
+Beyond resolving the invariance-equivariance trade-off, seq-JEPA further benefits from moving away from the two-view paradigm in SSL.
+We show that it performs well on tasks requiring integration over sequences of observations.
+In one scenario, inspired by embodied vision in primates, our model learns image representations without augmentations or masking, solely by predicting across simulated eye movements (saccades).
+In another setting, it performs path integration over sequences of actions—such as eye movements or 3D object rotations in 3DIEBench (Garrido et al., 2023).
+Our key contributions are as follows:
+  • We introduce seq-JEPA, a self-supervised world model that learns architecturally distinct equivariant and invariant representations through sequential prediction over action-observation pairs, without requiring explicit equivariance losses or dual predictors.
+  • We empirically validate that seq-JEPA matches or outperforms existing invariant and equivariant SSL methods across tasks requiring either representational property.
+  • We demonstrate that seq-JEPA naturally supports tasks that involve sequential integration of observations, such as predictive learning across saccades and path integration over action sequences.
 2 Method
 2.1 Invariant and equivariant representations
-Before describing the architecture and training procedure of seq-JEPA, we provide the nomenclature of invariant and equivariant representation learning in SSL (Dangovski et al., 2022; Devillers & Lefort, 2022).
-Consider $\mathcal{T}$ as a distribution of possible data transformations that is parametrized by a vector $t$.
-These transformations can be any type of hand-crafted data augmentation, such as geometric transformations or masking, that transform an original data point $x$ to different views.
-For example, we denote $x_{1}$ and $x_{2}$ as views generated from $x$ via two transformations $t_{1}$ and $t_{2}$ sampled from $\mathcal{T}$.
+Before presenting our architecture and training procedure, we briefly define invariance and equivariance in the context of SSL (Dangovski et al., 2022; Devillers and Lefort, 2022).
+Let $\mathcal{T}$ denote a distribution over transformations, parameterized by a vector $t$.
+These transformations—such as augmentations or masking—can be used to generate multiple views from a single image $x$.
+Let $x_{1}$ and $x_{2}$ be two such views, produced by applying transformations $t_{1}$ and $t_{2}$ sampled from $\mathcal{T}$.
+Additionally, let $a$ denote the relative transformation that maps $x_{1}$ to $x_{2}$.
 Additionally, we denote $a$ as a transformation that transforms $x_{1}$ to $x_{2}$.
-Here, we distinguish between $t$ and $a$, and call $t$ a transformation and $a$ an action, i.e., a relative transformation that transforms $x_{1}$ to $x_{2}$.
-Furthermore, we denote $f$ as an encoder function that projects the data space into a latent space.
-Given a transformation $t$, the encoder and its representations are said to be equivariant to $t$ if
-$\forall{t\in\mathcal{T}},~{}\exists{\,u_{t}}\;\mathrm{s.t.}\qquad f(t(x))=u_{t}(f(x)).$ (1)
-We can also define equivariance given the action $a$ and two views $x_{1}$ and $x_{2}$ as
-$\forall{a\in\mathcal{T}},~{}\exists{\,u_{a}}\;\mathrm{s.t.}\qquad f(x_{2})=u_{a}(f(x_{1})),$ (2)
-where $u_{t}$ and $u_{a}$ are functions that produce displacements in latent space corresponding to their associated transformation/action.
-As a special case of equivariance, the encoder and its representations are said to be invariant to $t$ if $u_{t}$ and $u_{a}$ are identity functions, i.e., $f(t(x))=f(x)$ or $f(x_{2})=f(x_{1})$.
+We distinguish between $t$ (an individual transformation) and $a$ (an action), where the latter reflects the change from one view to another.
+Let $f$ be an encoder that maps inputs to a latent space.
+We say that $f$ is equivariant to $t$ if:
+$\forall{t\in\mathcal{T}},~{}\exists{,u_{t}};\mathrm{s.t.}\qquad f(t(x))=u_{t}(f(x)),$ (1)
+where $u_{t}$ is a transformation in latent space corresponding to $t$. Equivariance can similarly be defined in terms of relative transformations (actions):
+$\forall{a\in\mathcal{T}},~{}\exists{,u_{a}};\mathrm{s.t.}\qquad f(x_{2})=u_{a}(f(x_{1})).$ (2)
+As a special case, if $u_{t}$ and $u_{a}$ are identity functions, then $f$ is invariant to the transformation: $f(t(x))=f(x)$ or $f(x_{2})=f(x_{1})$.
 2.2 Architecture
-The overall architecture of seq-JEPA is depicted in Figure 2.
-Denote $\{x_{i}\}_{i=1}^{M+1}$ as a sequence of views generated from a data sample $x$ using transformations $\{t_{i}\}_{i=1}^{M+1}$.
-The corresponding actions (relative transformations), whose embeddings are learned via a linear layer are denoted as $\{a_{i}\}_{i=1}^{M}$.
-A backbone encoder, $f_{\theta}$ encodes the first $M$ views and produces representations $\{z_{i}\}_{i=1}^{M}$.
-These representations, except for $z_{M}$, are concatenated with their corresponding action embeddings and fed to a sequence model, e.g., a transformer encoder, $g_{\phi}$ (no MLP projector is used after the encoder).
-This transformer module is tasked with aggregating the sequence of representations of action-observation pairs to produce an aggregate representation.
-We use a learnable token which we call [AGG] to learn the aggregate representation (this token resembles a [CLS] token in vision transformers (Dosovitskiy et al., 2020)).
-The output corresponding to [AGG] is therefore:
-$z_{AGG}=g_{\phi}((z_{1},a_{1}),(z_{2},a_{2})...,(z_{M-1},a_{M-1}),z_{M})$ (3)
-The representation $z_{AGG}$ is subsequently concatenated with the last action embedding $a_{M}$ which corresponds to the relative transformation between $x_{M}$ and $x_{M+1}$, and is fed to an MLP predictor module $h_{\psi}$ to predict the representation of $x_{M+1}$, i.e., $\hat{z}_{M+1}=h_{\psi}(z_{AGG},a_{M})$.
-The last view $x_{M+1}$, is itself encoded using a target encoder whose weights are an exponential moving average (EMA) of the encoder $f_{\theta}$ used to encode the sequence of observations.
-The representation $z_{M+1}$ is passed through an stop-gradient (sg) to prevent representational collapse.
-Our objective is then to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$, which we achieve by minimizing the following prediction loss,
-$\mathcal{L}_{seq-JEPA}=1-\frac{{\hat{z}_{M+1}}}{\left\lVert{{\hat{z}_{M+1}}}\right\rVert_{2}}\cdot\frac{{\texttt{sg}(z_{M+1}})}{\left\lVert{{\texttt{sg}(z_{M+1}})}\right\rVert_{2}}.$ (4)
-No additional equivariance predictor or loss function is used when training the framework.
-Figure 2: Overview of the proposed architecture; seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn separate representations for equivariance- and invariance-related tasks.
+Figure 1 presents the overall architecture of seq-JEPA.
+Let $\{x_{i}\}_{i=1}^{M+1}$ be a sequence of views generated from a sample $x$ via transformations $\{t_{i}\}_{i=1}^{M+1}$.
+The relative transformations (actions) $\{a_{i}\}_{i=1}^{M}$ are defined as $a_{i}\triangleq\Delta t_{i,i+1}$, i.e., the transformation mapping $x_{i}$ to $x_{i+1}$.
+In our default setting, we use a learnable linear projector to encode these actions.
+A backbone encoder, $f$ encodes the first $M$ views, producing representations $\{z_{i}\}_{i=1}^{M}$.
+Except for $z_{M}$, each $z_{i}$ is concatenated with its corresponding action embedding and passed to a transformer encoder $g$ (no MLP projector is used after the encoder), which aggregates the sequence of action-observation pairs.
+The transformer uses a learnable [AGG] token (analogous to the [CLS] token in ViTs (Dosovitskiy et al., 2020)) to generate the aggregate representation:
+$z_{AGG}=g((z_{1},a_{1}),(z_{2},a_{2})...,(z_{M-1},a_{M-1}),z_{M})$ (3)
+This aggregate representation $z_{AGG}$ is then concatenated with the final action embedding $a_{M}$ (corresponding to the transformation from $x_{M}$ to $x_{M+1}$), and passed to an MLP predictor $h$ to predict the representation of $x_{M+1}$:
+$\hat{z}_{M+1}=h(z_{AGG},a_{M}).$ (4)
+The ground truth $z_{M+1}$ is computed using a target encoder—an exponential moving average (EMA) of $f$.
+The target representation is passed through a stop-gradient operator (sg) to avoid representational collapse.
+The training objective is to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$:
+$\mathcal{L}_{\mathrm{seq-JEPA}}=1-\frac{\hat{z}_{M+1}}{\left\lVert{\hat{z}_{M+1}}\right\rVert_{2}}\cdot\frac{\texttt{sg}(z_{M+1})}{\left\lVert{\texttt{sg}(z_{M+1})}\right\rVert_{2}}.$ (5)
+No additional loss terms or equivariance-specific predictors are used during training.
+Figure 1: seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn architecturally distinct invariant and equivariant representations.
+Figure 2: a.
+Transformations and observations used for training b.
+In predictive learning across saccades, image saliencies and inhibition of return help create more informative and less redundant patch sequences.
 2.3 Action and observation sets
-In order to evaluate our model and measure its ability to generalize to different types of transformations, we consider three distinct sets of transformation and their corresponding action-observation pairs (Figure 3a).
-Descriptions of each transformation set are provided below, with additional details available in Section A.1.
+To evaluate generalization across transformation types, we define three sets of action-observation pairs (Figure 2). See Appendix A.1 for details of each setup.
 3D Invariant Equivariant Benchmark (3DIEBench).
-We use the 3DIEBench dataset (Garrido et al., 2023) as our first evaluation benchmark.
-This dataset is specifically designed for measuring equivariance and provides 3D object renderings with different object rotations and floor hue and lighting angles and their corresponding parameters.
-We specifically study equivariance to the 3D rotation group ($SO(3)$) and further to floor and light hue in this benchmark.
-Hand-crafted data augmentations.
-In this setting, observations are transformed image views generated by hand-crafted augmentations that are typically used in the two-view paradigm of SSL.
-The actions are relative augmentation parameters between two views.
-We measure equivariance to crop, color jitter, and blur transformations.
-As benchmarks for this setting, we use CIFAR100 and Tiny ImageNet datasets.
-To generate augmented views, we follow EquiMod’s augmentation strategy (Devillers & Lefort, 2022).
-For details on augmentations and the corresponding parameterization, see Section A.1.
-Predictive learning across saccades.
-Going beyond transformations such as augmentations and rotation, we explore other ways of leveraging sequential observations to learn visual representations.
-Specifically, we designed a procedure to learn image representations without using any hand-crafted augmentations with a similar flavor to I-JEPA (Assran et al., 2023) but without dependence on masking.
-Our approach is enabled by seq-JEPA’s architecture and its ability to produce an aggregate representations from a sequence of observations.
-In this setting, given a full-size high-resolution image, we extract a number of smaller patches and use them as a sequence of partial observations.
-In our experiments with the STL-10 dataset, we use 32 by 32 patches, i.e.
-one-ninth of the original image.
-The action in this setup corresponds to the relative position between the center of two subsequent patches, i.e.
-a simulated act of rapid eye movements (also known as saccades) from one point of the image to another.
-Although it is possible to use random positions as patch centers (fixations) for the sequence, we show that adopting two techniques inspired by embodied vision in primates improves the quality of the aggregate representation and its downstream classification performance (Figure 3b).
-The first technique utilizes image saliencies (Itti et al., 1998; Li, 2002; Zhaoping, 2014) to sample fixations.
-To generate saliency maps, we use a pre-trained saliency prediction model, DeepGaze IIE (Linardos et al., 2021).
-Given a dataset, we pre-extract and store the saliencies without any additional overhead during training.
-These saliency maps can be interpreted as how likely humans are to center their gaze (fixate) on a given pixel of the image.
-Therefore, sampling fixations from the probability distribution corresponding to saliency maps provides more informative patches, yet similar to uniform sampling, is prone to a high degree of redundancy and overlap between patches.
-To remedy this, i.e.
-minimize overlap among patches and avoid future fixation that revisit previously attended areas, we use another technique inspired by biological vision known as inhibition of return (IoR) (Posner et al., 1985).
-To implement IoR, after sampling a fixation from saliency maps, we zero out the sampling probabilities of an area with a diameter equal to the patch width surrounding the previous fixations before sampling a subsequent fixation.
-In this scenario, we are interested to see whether encoder representations become sensitive to fixation positions (spatial awareness), and therefore equivariance is measured with respect to fixation coordinates.
-Figure 3: a.
-Types of transformations and observations used for training; seq-JEPA works with a variety of action-observation pairs such as 3D object rotations, hand-crafted augmentations, and simulated saccades.
-b.
-For predictive learning across saccades, image saliencies and inhibition of return help create more informative and less redundant patch sequences.
-3 Related Work
-Joint-embedding predictive architectures.
-In JEPAs (LeCun, 2022; Dawid & LeCun, 2024), a predictor asymmetry is introduced in one of the two branches that traditionally encode transformed views in SSL.
-The representation learning task is then framed as a world modeling objective (Ha & Schmidhuber, 2018; Hafner et al., 2019), i.e., the predictor is tasked with predicting the consequence of a data transformation (action) applied to the first view that results in the second view.
-The predictor is conditioned on transformation parameters such as positional masks to predict the latent representation of target blocks corresponding to the masks as in I-JEPA (Assran et al., 2023), or additionally on parameters corresponding to hand-crafted image augmentations as in Image World Model (IWM) (Garrido et al., 2024).
-Recent works have also explored JEPAs for physical understanding (Garrido et al., 2025) and reward-free offline planning (Sobal et al., 2025).
-Non-generative world models.
-Non-generative world models predict the consequence of an action or transformation in latent space.
-A line of work considers such world models in contrastive SSL by predicting the representation of a transformed view from another via a contrastive objective and conditioned on the past or context (van den Oord et al., 2019; Gupta et al., 2024).
-The family of JEPAs defer the task of world modeling to a predictor module that is conditioned on the action (Assran et al., 2023; Garrido et al., 2024; Sobal et al., 2025).
-Non-generative world models have also garnered attention in model-based reinforcement learning (RL) and planning to improve sample efficiency (Schwarzer et al., 2021), to learn environment dynamics (Sobal et al., 2025), or to create an intrinsic novelty-based reward to improve exploration in self-predictive world models (Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023).
-A prominent example of such world models is BYOL-Explore (Guo et al., 2022), in which a recurrent cell follows the encoder and is conditioned on the current action in a planning task, followed by a predictor to capture the environment’s transition dynamics.
-Equivariant representation learning.
-Many equivariant SSL methods extend invariant methods by introducing an additional equivariance objective.
-This objective can be direct prediction of parameters corresponding to augmented views (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024), or additional transformations such as rotation (Dangovski et al., 2022).
-EquiMod (Devillers & Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space in addition to their invariant objective.
-Xiao et al.
-(2021) use contrastive learning to train multiple projection heads for equivariance to each augmentation, where the views corresponding to the desired augmentation are considered as negative samples.
-In SEN (Park et al., 2022), similar to EquiMod, a predictor is used to predict the effect of transformations in latent space, yet without any invariance objective.
-ContextSSL (Gupta et al., 2024) learns representations that are equivariant only to the recent transformations in the current context by conditioning representations on both the current action and the past via a memory module, and further using a transformation prediction objective to avoid collapse to invariance.
-Another line of work including Shakerinava et al.
-(2022), CARE (Gupta et al., 2023), and Yerxa et al.
-(2024) do not assume access to transformation parameters yet require two pairs of views to which the same transformation is applied, and an equivariance loss function is minimized to induce equivariance in representations.
-Generative SSL methods based on masked image modeling (He et al., 2022; Bao et al., 2021; Xie et al., 2022) can become equivariant to position due to the use of positional embeddings.
-Finally, action-conditioned JEPAs condition the predictor on positional mask tokens (Assran et al., 2023) and augmentation parameters (Garrido et al., 2024) to induce equivariance in encoder representations without any additional equivariance objective.
-Relationship to existing literature.
-Our framework belongs to the family of JEPAs and is a non-generative world model.
-We show that via architectural inductive biases, JEPAs can be leveraged to resolve the invariance-equivariance trade-off in SSL and enable processing of sequential action-observation pairs that is required to solve specific perception tasks.
-Unlike the majority of equivariant SSL methods discussed above, seq-JEPA does not use an equivariance predictor or loss function, nor does it require two pairs of views to which the same transformation is applied.
-We further highlight the differences between our framework and a recent equivariant method, ContextSSL.
-Specifically, seq-JEPA can be distinguished from ContextSSL on two aspects.
-First, unlike seq-JEPA, ContextSSL operates within the two-view SSL paradigm but uses a decoder-only transformer instead of the usual MLP projector so that its output embeddings are conditioned on the past.
-Second, while ContextSSL’s main goal is adapting equivariance to different transformations by alternating transformations in training data, our goal is to learn separate invariant and equivariant representations with respect to one or more transformations specified, and solve downstream tasks that require aggregating a sequence of action-observation pairs.
+We first evaluate on the 3DIEBench dataset (Garrido et al., 2023), designed for equivariance analysis.
+It includes 3D object renderings with variations in rotation, floor hue, and lighting.
+We primarily study equivariance to $SO(3)$ rotations and secondarily to appearance factors of floor and light hue.
+Hand-Crafted Augmentations.
+In this setting, we use transformed views generated via common SSL augmentations (e.g., crop, color jitter, blur), where actions correspond to relative augmentation parameters.
+We use CIFAR100 and Tiny ImageNet, and follow EquiMod’s augmentation protocol (Devillers and Lefort, 2022).
+Predictive Learning Across Saccades (PLS).
+Going beyond conventional transformations such as augmentations or 3D rotations, we show that seq-JEPA can learn visual representations from a sequence of partial observations thanks to architectural inductive biases—without relying on any hand-crafted augmentations.
+Our PLS has a similar flavor to I-JEPA (Assran et al., 2023) but does not require engineered masking strategies.
+In PLS, we train seq-JEPA on sequences of patches extracted from high-resolution images.
+For instance, with STL-10, we use $32\times 32$ patches to form the observation sequence.
+In this setting, actions correspond to the relative positions between patch centers, simulating saccadic eye movements and inducing 2-D positional equivariance to representations.
+To select fixation points, we adopt two biologically inspired techniques that increase informativeness, reduce redundancy, and improve the downstream utility of the aggregate representation (Figure 2):
+  • Saliency-Based Fixation Sampling.
+Using DeepGaze IIE (Linardos et al., 2021), we extract saliency maps for each image and use them to probabilistically sample fixation points (Itti et al., 1998; Li, 2002; Zhaoping, 2014).
+The maps are pre-computed and introduce no training overhead.
+  • Inhibition of Return (IoR).
+To reduce spatial overlap between patches and emulate natural exploration (Posner et al., 1985), we implement IoR by zeroing out the sampling probability of areas surrounding previously sampled fixations.
 3 Experimental Setup
 3.1 Compared methods and baselines
-We compare seq-JEPA’s performance with both invariant and equivariant SSL methods.
-For invariant methods, we compare with SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022).
-For equivariant methods, we compare with SEN (Park et al., 2022), EquiMod (Devillers & Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024).
-Additionally, we consider another equivariant baseline which we call Conditional BYOL, where BYOL’s predictor is conditioned on the action that transforms the view of the target branch to the view of the online branch, with the aim that transformation information is also encoded in representations.
-This architecture is equivalent to a two-view version of seq-JEPA with a sequence length of one during both training and inference, and with no sequence aggregator on top of representations.
-Furthermore, as a baseline for predictive learning across saccades, we consider a model which we call Conv-JEPA.
-This baseline has the same architecture as Conditional BYOL, and to train it, we consider the same sequence of patches that seq-JEPA uses (sampled from image saliencies), predict the representation of the last patch from each individual patch in the sequence, and sum their prediction losses.
-For architectural details of each baseline, see Section A.3.
+We compare seq-JEPA against both invariant and equivariant SSL baselines.
+Invariant methods include SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022).
+Equivariant methods include SEN (Park et al., 2022), EquiMod (Devillers and Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024).
+For all baselines, architectural details are given in Appendix A.3.
+We also evaluate two hybrid baselines based on our architecture:
+  • Conditional BYOL.
+A two-view version of seq-JEPA with no sequence aggregator, where BYOL’s predictor is conditioned on the relative transformation between target and online views.
+This encourages representations to encode transformation information.
+  • Conv-JEPA.
+A baseline for the saccades setting.
+It uses the same sequence of saliency-sampled patches as seq-JEPA and predicts the final patch’s representation from each earlier patch individually.
+These losses are summed across the pairs before backpropagation.
 3.2 Training protocol
-For all methods, we use a ResNet-18 (He et al., 2016) (not pretrained) as the backbone encoder.
-For experiments that use CIFAR100 and low-resolution STL-10 patches in predictive learning across saccades, we use the CIFAR variant of ResNet-18.
-For action conditioning, similar to EquiMod (Devillers & Lefort, 2022), we use a learnable linear projection to learn action embeddings.
-In our default setting for all experiments with action conditioning, the action projector is 128-dimensional.
-For the sequence aggregator, we use a small transformer encoder (Vaswani et al., 2017) with three layers, four attention heads, and post-normalization.
-For the predictor, we use an MLP with a hidden layer of 1024 dimensions and ReLU activation.
-In baseline comparisons, in order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train two variants of baselines that typically employ an MLP projector: one maintaining the original MLP projector and another where the projector is replaced with our transformer encoder.
-We report the higher performance among the two models.
-For a fair comparison, all models are trained for a fixed batch size of 512 and the same number of epochs.
-For 3DIEBench experiments, all models are trained for 1000 epochs, while for other datasets, they are trained for 2000 epochs.
-For models that use a transformer as projector, including seq-JEPA, we use the AdamW optimizer with default $\beta_{1}$ and $\beta_{2}$, a weight decay of $0.001$, and a learning rate of $4\times 10^{-4}$ with a linear warmup for $20$ epochs starting from $10^{-5}$ followed by a cosine decay back to $10^{-5}$.
-For models that use an MLP projector, we use the Adam optimizer with a learning rate of $10^{-3}$, default $\beta_{1}$ and $\beta_{2}$, and no weight decay.
+All models use ResNet-18 (He et al., 2016) as the backbone encoder.
+For action conditioning, we use a learnable linear projection to learn action embeddings (default action embedding is 128-d).
+The sequence aggregator in seq-JEPA is a lightweight transformer encoder (Vaswani et al., 2017) with three layers and four attention heads.
+The predictor is a 2-layer MLP with 1024 hidden units and ReLU activation.
+In order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train baselines that typically use MLP projectors in two variants: (1) with original MLP projector; and (2) with the MLP replaced by a transformer encoder and a sequence length of one.
+We report the better-performing variant to ensure a fair comparison.
+All models are trained from scratch with a batch size of 512.
+We use 1000 epochs for 3DIEBench and 2000 epochs for other datasets to obtain asymptotic performance.
+We use AdamW for models with transformer projectors (including seq-JEPA) due to its stability and improved regularization in transformer training.
+For ConvNet-only models with MLP heads, we use the Adam optimizer.
+Full hyperparameters are detailed in Appendix A.
 3.3 Evaluation metrics and protocol
-To measure equivariance with respect to a specific transformation, we perform action decoding from encoder representations as in Garrido et al.
-(2023).
-Specifically, given two transformed views, we extract representations from the frozen encoder, concatenate them, and train a regression head on top to predict the relative transformation (action) between the two.
-We report the $R^{2}$ score of this regressor on the test set as the equivariance metric.
-Although not always a perfect measure of invariance (Wang et al., 2024), we report top-1 classification accuracy of a linear probe as an approximate metric for representational invariance.
-For all baselines, we report the accuracy based on encoder representations.
-For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 2).
-When reporting accuracy based on $z_{AGG}$, we report the number of observation views (context lengths) used during both training and inference.
-For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 2) for all seq-JEPA models in Section B.4.
-Details on training the evaluation probes are given in Section A.2.
+To assess equivariance, we follow the protocol of Garrido et al.
+(2023) and train a regressor on frozen encoder representations to predict the relative transformation (action) between two views.
+We report the $R^{2}$ score on the test set.
+As a proxy measure of invariance, we use top-1 classification accuracy of a linear probe on top of frozen representations.
+For all baselines, probes are trained on encoder outputs.
+For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 1) and report the number of observation views used during training and inference.
+For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 1) for seq-JEPA models in Appendix B.4.
+Training details of evaluation heads are given in Appendix A.2.
 4 Results
@@
 We use the 3DIEBench benchmark to quantitatively compare invariant and equivariant representations in seq-JEPA with baseline methods.
-This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object category classification.
-Table 1 provides a summary of our evaluation results on the 3DIEBench dataset where equivariant methods have been conditioned on rotation.
-In addition to the relative rotation between two views, in the last column, we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view.
-For seq-JEPA, we trained models with training sequence lengths between one to four (denoted by $M_{tr}$ in the table) and measure the linear classification performance with different inference lengths (denoted by $M_{val}$) on top of the aggregate representation.
-Among the invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance.
-Among equivariant baselines, while EquiMod and SEN offer even better classification accuracy compared to invariant methods, in terms of equivariance, they lag behind SIE and ContextSSL which provide the best rotation prediction performance due to their specialized equivariance predictors and loss functions.
-Even though both SIE and ContextSSL representations become highly sensitive to rotation, they do not fare well in linear probe classification.
-Seq-JEPA is able to achieve a rotation $R^{2}$ on a par with the best performing equivariant method while outperforming all methods in classification thanks to sequence aggregation during training and inference with a margin that increases with the inference sequence length.
-Finally, when action conditioning is ablated (last row in the table), seq-JEPA loses its equivariant properties, yet maintains a strong classification performance.
-For additional results on 3DIEBench including models with different training and inference sequence lengths and models conditioned on rotation and color, see Sections B.1 to B.3.
-Table 1: Quantitative evaluation of learned representations on 3DIEBench for rotation prediction (equivariance) and linear probe classification (invariance).
+This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object classification.
+Table 3 provides a summary of our evaluation on the 3DIEBench where equivariant methods have been conditioned on rotation.
+In addition to the relative rotation between two views, in the last column we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view.
+For seq-JEPA, we trained models with varying training sequence lengths (denoted by $M_{tr}$ in the table) and measure the linear classification performance on top of aggregate representations with different inference lengths (denoted by $M_{val}$).
+Among invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance.
+Among equivariant baselines, SIE and ContextSSL yield strong rotation prediction performance due to their specialized equivariance predictors and loss functions, but underperform in classification.
+EquiMod and SEN offer better classification performance, yet compromise equivariance.
+In contrast, seq-JEPA achieves strong performance in both, matching the best rotation $R^{2}$ scores while exceeding baselines in classification, with gains increasing with inference sequence length $M_{val}$.
+Ablating action conditioning leads to a sharp drop in equivariance but retains classification accuracy, confirming our hypothesis that action-conditioned sequential aggregation enables a segregated invariant-equivariant representation structure.
+For additional results on 3DIEBench including models with varying training and inference sequence lengths and models conditioned on both rotation and color, see Appendices B.1 to B.3.
+Figure 3: Top-1 linear classification (invariance) vs.
+rotation prediction (equivariance) performance on 3DIEBench.
+seq-JEPA learns good representations for both tasks.
+Subscripts indicate training and inference sequence lengths.
+Table 1: Evaluation on 3DIEBench for linear probe classification (invariance) and rotation prediction (equivariance).
 Equivariant models and seq-JEPA are conditioned on rotation.
-For seq-JEPA, the training sequence length is denoted by $M_{tr}$ and probing on top of $z_{AGG}$ is done with an inference length $M_{val}$.
-Method Classification (top-1) Relative Rotation ($R^{2}$) Individual Rotation ($R^{2}$)
+For seq-JEPA, training and inference sequence lengths are denoted by $M_{tr}$ and $M_{val}$.
+Method Top-1 Acc. (%) Rel. Rot. ($R^{2}$) Indiv. Rot. ($R^{2}$)
 Invariant
@@
 Conditional BYOL 82.61 0.31 0.47
-ContextSSL ($c=126$) 80.4 0.74 0.78
+ContextSSL ($c=126$) 80.40 0.74 0.78
 Invariant-Equivariant
-seq-JEPA ($M_{tr}=1$, $M_{val}=1$) 84.08 0.65 0.69
-seq-JEPA ($M_{tr}=1$, $M_{val}=3$) 85.31 0.65 0.69
-seq-JEPA ($M_{tr}=3$, $M_{val}=3$) 86.14 0.71 0.74
-seq-JEPA ($M_{tr}=3$, $M_{val}=5$) 87.41 0.71 0.74
-seq-JEPA (w/o act cond, $M_{tr,val}=3$) 86.05 0.29 0.37
+seq-JEPA (1, 1) 84.08 0.65 0.69
+seq-JEPA (1, 3) 85.31 0.65 0.69
+seq-JEPA (3, 3) 86.14 0.71 0.74
+seq-JEPA (3, 5) 87.41 0.71 0.74
+seq-JEPA (no act cond) 86.05 0.29 0.37
 4.2 Qualitative evaluation on 3DIEBench
-In order to qualitatively evaluate the representations learned by seq-JEPA and shed some light on its invariance-equivariance properties, we conduct two different visualizations.
-First, to see equivariance in representational space, we retrieve the three nearest representations of a query image from the validation set of 3DIEBench (Figure 4).
-All methods successfully retrieve the correct object category, while only seq-JEPA and SIE, which demonstrate high quantitative equivariance performance (Table 1), maintain rotation equivariance across all three nearest representations.
-Figure 4: Retrieval of nearest representations; given a query image, we extract the three nearest encoder representations in the validation set of 3DIEBench.
-The retrieved views of models with the highest quantitative rotation equivariance performance maintain the rotation of the query image.
-In the second visualization, we plot the 2D UMAP projections of seq-JEPA’s encoder and aggregate representations (Figure 5) for the model trained conditioned on rotation.
-The left panel of of Figure 5 shows encoder representations colored by class label, while the middle panel displays the same encoder representations colored by rotation angle.
-The smooth color gradation across the map within each class cluster in this middle panel suggests that the encoder captures rotation angle as a continuous factor, implying equivariance to rotation (e.g., the red class in the bottom-right corner of the right panel and the corresponding part in the middle panel).
+To visualize equivariance in representational space, we retrieve the three nearest representations of a query image from the validation set of 3DIEBench (Figure 5).
+While all models retrieve the correct object category, only seq-JEPA and SIE consistently preserve rotation across all retrieved views, consistent with their high $R^{2}$ scores.
+Next, we projected encoder and aggregate representations using 2D UMAP (Figure 5).
+The left panel shows encoder representations colored by class label, while the middle panel displays the same encoder representations colored by rotation angle.
+The smooth color gradation across the map within each class cluster in the middle panel suggests that the encoder captures rotation angle as a continuous factor, implying equivariance to rotation (e.g., the red class in the bottom-right corner of the right panel and the corresponding part in the middle panel).
 The right panel shows aggregate representations colored by class label.
 Comparing the class-colored plots (left and right panels), we observe that both encoder and aggregate representations contain class information.
-However, when we aggregate multiple views of a sample, some of the intra-class variability (resulting from transformations such as rotation) is eliminated, causing each class’s representational cluster to become more homogeneous.
+However, when we aggregate multiple views of a sample, some of the intra-class variability (resulting from transformations such as rotation) is eliminated, causing each class’ representational cluster to become more homogeneous.
 This aggregation procedure likely reduces variation due to rotation and makes the representations more invariant, resulting in decreased intra-class spread and increased inter-class distance.
-We create a similar UMAP visualization for seq-JEPA with ablated rotation conditioning in Section B.5 which highlights the role of action conditioning in achieving equivariance to rotation.
-Figure 5: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations.
-The model is trained on 3DIEBench and conditioned on rotation with $M_{tr}=3$ and $M_{val}=5$.
-Encoder representations for each view observation, color-coded by class category (left) and rotation angle (middle).
-The aggregate token representation for $M_{val}=5$ observations during inference, color-coded by class category (right).
-5.3 Evaluation of models trained using hand-crafted augmentations
-We next quantitatively assess invariance and equivariance for the second set of transformations, i.e.
-hand-crafted augmentations.
-Table 2 provides a summary of our evaluation on models trained on CIFAR100 and Tiny ImageNet.
-For each equivariant method and seq-JEPA, we trained four different variations.
-In three variations, the model is conditioned on only one of the augmentations among cropping, color jittering, and blurring, and in the last variation, on all three augmentations.
-Considering equivariance to augmentations, seq-JEPA outperforms invariant and equivariant baselines by a large margin.
+We create a similar UMAP visualization for seq-JEPA with ablated rotation conditioning in Appendix B.5 to highlight the role of action conditioning in achieving equivariance to rotation.
+Figure 4: Retrieval of nearest representations; given a query image, we extract the three nearest encoder representations in the validation set of 3DIEBench.
+The retrieved views of models with the highest quantitative rotation equivariance performance maintain the rotation of the query image across all retrieved views.
+Figure 5: 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations trained on 3DIEBench and conditioned on rotation with $M_{tr}=3$ and $M_{val}=5$.
+Encoder representations for each view observation, color-coded by class (left) and rotation angle (middle).
+Aggregate representation for $M_{val}=5$, color-coded by class (right).
+4.3 Evaluation with Hand-Crafted Augmentations
+We assess invariance and equivariance under hand-crafted augmentations by training on CIFAR100 and Tiny ImageNet (Table 2).
+Models with action conditioning are trained conditioned on crop, color jitter, blur, or all three (indicated in the first column of the table).
+seq-JEPA consistently achieves higher equivariance than both invariant and equivariant baselines across all transformations.
 Notably, except for the model trained on CIFAR-100 and conditioned on blur, the best equivariance performance for a given augmentation is achieved when the model is specialized and conditioned only on that augmentation.
-Furthermore, ablating actions (last row in the table) causes seq-JEPA to achieve a lower equivariance performance across transformations compared to action-conditioned models.
-In terms of linear probe classification accuracy on top of aggregate representations, our framework is competitive with both invariant and equivariant families, yet does not outperform all of them.
-For additional results with different training and inference sequence lengths for seq-JEPA models trained with augmented views, see Section B.3.
-Table 2: Quantitative evaluation of learned representations with hand-crafted augmentations on CIFAR100 and Tiny ImageNet; equivariance is measured by predicting relative transformation parameters associated with crop, color jitter, or blur augmentations.
-For seq-JEPA, the training sequence length is denoted by $M_{tr}$ and probing is done with an inference length ($M_{val}$) of five.
+Furthermore, ablating actions (last row in the table) causes seq-JEPA to lose its equivariance across transformations compared to action-conditioned models.
+Overall, our model outperforms both invariant and equivariant families in terms of equivariance, while being competitive in terms of classification performance.
+For additional results with varying training and inference sequence lengths, see Appendix B.3.
+Table 2: Evaluation with hand-crafted augmentations on CIFAR100 and Tiny ImageNet; equivariance is measured by predicting relative transformation parameters associated with crop, color jitter, or blur augmentations.
+For all seq-JEPA models, $M_{val}=5$.
 CIFAR100 Tiny ImageNet
@@
 - seq-JEPA ($M_{tr}=2$) 58.37 0.64 0.14 0.16 35.97 0.52 0.18 0.47
-5.4 Predictive learning across saccades
-Quantitative evaluation.
-In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations alongside actions to learn visual representations.
-Table 3 summarizes our evaluation results for predictive learning across saccades on STL-10.
-While seq-JEPA only uses low-resolution patches with no augmentation, it achieves a classification performance comparable to SimCLR, an invariant model that is trained using strongly augmented image views with full resolution.
-Ablating action conditioning causes the accuracy on top of the aggregate representations to drop significantly, indicating that positional awareness is essential to forming semantic representations across simulated eye movements.
-Furthermore, comparison with the Conv-JEPA baseline highlights the importance of aggregating a sequence of patch representations in contrast to simply accumulating prediction errors between pairs of patches.
-Additionally, the last two ablation results in the table show that saliency-driven sampling of patches alongside IoR play a key role in sampling informative and non-overlapping patches and achieving high-quality aggregate representations.
-Interestingly, while random uniform patch sampling negatively impacts classification accuracy, it results in the highest positional equivariance, as the model samples patches and corresponding saccade actions from across the entire image and not just the salient regions.
-Table 3: Quantitative evaluation of predictive learning across saccades on STL-10.
-Equivariance is measured with respect to the position of a fixated patch on the image.
-Seq-JEPA models use only low-resolution patches with no augmentation during training and inference.
-Unless stated otherwise, training and inference lengths ($M_{tr}$ and $M_{val}$) are set to four for all seq-JEPA variants.
-Conditioning Classification (top-1) Position ($R^{2}$)
+4.4 Predictive Learning across Saccades and Path Integration
+In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations to learn visual representations in a predictive manner.
+Table 6 shows that seq-JEPA reaches 83.44% top-1 accuracy on STL-10, comparable to SimCLR (85.23%) trained with full-resolution images and strong augmentations.
+This gap narrows further when increasing the inference sequence length from $M_{val}=4$ to $M_{val}=6$.
+Ablating action conditioning causes the accuracy on top of the aggregate representations to drop sharply, indicating that 2-D positional awareness is essential to forming semantic representations across simulated eye movements.
+Compared to Conv-JEPA, which accumulates prediction losses pairwise, seq-JEPA performs better in classification, highlighting the importance of sequence aggregation when dealing with partial observations in SSL.
+Further ablations show that saliency-driven sampling and IoR are critical for forming informative, non-overlapping patch sequences and subsequently a high-quality aggregate representation.
+Interestingly, while random uniform patch sampling negatively impacts classification accuracy due to lower semantic content, it results in the highest positional equivariance as the model samples patches and corresponding saccade actions from a more diverse set of positions across the entire image, not just the salient regions.
+UMAP projections for PLS with and without action conditioning in Appendix B.5 further underscore the role of action conditioning in enabling positional equivariance.
+Path Integration.
+In the context of eye movement-driven observations and in general sequential observations, another ability that naturally arises is path integration (McNaughton et al., 2006)—predicting the cumulative transformation/action from a sequence of actions.
+We evaluate this task in both eye movements in PLS (visual path integration) and object rotations in 3DIEBench (angular path integration).
+As shown in Figure 6, seq-JEPA demonstrates strong performance in both settings, with performance degrading gracefully as sequence length increases.
+Ablating action conditioning causes path integration to fail, whereas ablating the visual stream has only a minor impact—highlighting that action information is the dominant signal for this task.
+Full details of the path integration setup are provided in Appendix B.6.
+Table 3: Evaluation of predictive learning across saccades on STL-10. Equivariance is measured with respect to fixation coordinates. Unless stated otherwise, $M_{tr}=M_{val}=4$.
+Conditioning Top-1 Acc. (%) Position ($R^{2}$)
 Invariant
-- SimCLR (augmentations) 85.23 -
+- SimCLR (augmentations) 85.23 -0.06
 Equivariant
@@
 position seq-JEPA ($M_{val}=6$) 84.12 0.80
-position seq-JEPA (no saliency) 79.85 0.88
-position seq-JEPA (no IoR) 77.97 0.85
-Qualitative evaluation.
-Figure 6 shows UMAP projections for seq-JEPA representations trained on STL-10 via predictive learning across saccades.
-As for rotation in Figure 5, equivariance to fixation position can be observed from the bottom panels that show UMAP projection of encoder representations color-coded by X and Y coordinates.
-Similar to rotation in 3DIEBench, we provide a similar UMAP visualization for seq-JEPA with ablated position conditioning in Section B.5 for predictive learning across saccades to highlight the role of action conditioning.
-Figure 6: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations.
-The model is trained on STL-10 via predictive learning across saccades and conditioned on fixation position with $M_{tr,val}=4$.
-Encoder representations for each view observation, color-coded by class category (top-left) and $X$ and $Y$ pixel coordinates (bottom panels).
-The aggregate token representation for $M_{val}=4$ observation patches during inference, color-coded by class category (top-right).
-5.5 Path integration across actions
-While an agent executes a sequence of actions in an environment, transitioning from an initial state to a final state, it should be capable of tracking its position by integrating its own actions.
-This is also a crucial cognitive ability that enables animals to estimate their current state in their habitat (McNaughton et al., 2006).
-Here, we evaluate whether pre-trained seq-JEPA is able to solve path integration.
-Given the sequence of observations $\{x_{i}\}_{i=1}^{M+1}$ generated from transformations $\{t_{i}\}_{i=1}^{M+1}$ and the corresponding action embeddings $\{a_{i}\}_{i=1}^{M}$, we define the task of path integration over the sequence of actions as predicting the relative action that would directly transform $x_{1}$ to $x_{M+1}$ (which we denote as $a_{1,M}$) given $z_{AGG}$ and $a_{M}$.
-In other words, given the aggregate representation of a sequence of action-observation pairs and the next action, we would like to predict the overall position change from the starting point ($x_{1}$) to the end point ($x_{M+1}$).
-We consider path integration for rotation angles in 3DIEBench and across eye movements with STL-10.
-For rotations, the task is integrating a series of object rotations from the first view to the last, i.e.
-angular path integration.
-For eye movements, the task is integrating the eye movements from the first saccade to the last, i.e.
-visual path integration.
-To measure path integration performance, we train a regression head on top of the concatenation of $z_{AGG}$ and $a_{M}$ to predict $a_{1,M}$ for a given inference sequence length.
-Figure 7 shows the results for both angular and visual path integration.
-The red curve corresponds to the performance of the original seq-JEPA.
-The blue curve corresponds to experiments in which the action embeddings are ablated (they are set to zero during inference for all views).
-The green curve corresponds to experiments in which the encoder (visual) representations are ablated (they are set to zero when being concatenated with actions before being fed to the transformer).
-As expected, path integration becomes more difficult as the number of observations increases (red curves).
-Ablating action conditioning (blue curves) results in failure of path integration.
-On the other hand, ablating the visual representations (green curves) results only in a small performance drop compared to the original model, indicating that action conditioning is the key factor enabling path integration.
+position seq-JEPA (w/o saliency & IoR) 79.85 0.88
+position seq-JEPA (w/o IoR) 77.97 0.85
 Figure 6: a) Visual path integration across eye movements, b) Angular path integration across object rotations (results over three random seeds)
-5.6 Role of action conditioning
-To better understand the importance of action conditioning in learning highly equivariant representations, we perform ablation experiments for seq-JEPA trained on 3DIEBench by changing the dimensionality of action embeddings and ablating actions from one of the two or both components of seq-JEPA, i.e.
-the transformer encoder and the MLP predictor.
-These ablation results are provided in Table 4.
-An important observation is that both predictor and transformer action conditioning are important in achieving the best equivariant performance, yet predictor action conditioning plays a more important role.
-Furthermore, learnable action embeddings play a key role in developing equivariance to rotation, although this equivariance shows little sensitivity to the dimensionality of the action embedding.
-Additionally, classification performance is slightly boosted in the presence of action conditioning, yet its role in this case is not as important compared to positional embeddings in predictive learning across saccades.
-Table 4: Ablation results on the effect of seq-JEPA’s action conditioning trained on 3DIEBench (equivariance to rotation); we investigate the effects of partial or complete action ablations and changing the dimensionality of action embeddings on linear classification and rotation prediction performance.
-All models have $M_{tr}=3$ and $M_{val}=5$ and results are reported over three randoms seeds.
-Variant Classification (top-1) Rotation ($R^{2}$)
+4.5 Ablations: Action Conditioning Is Key
+To better understand the mechanisms underlying seq-JEPA’s invariant-equivariant representation learning and role of action conditioning, we perform a set of ablation experiments on 3DIEBench.
+Specifically, we study: (i) the role of action conditioning in the transformer and predictor; (ii) the impact of learning action embeddings and their dimensionality.
+Table 7 summarizes our ablation results.
+Removing action conditioning entirely causes a significant drop in equivariance ($R^{2}$ from 0.71 to 0.29), although classification accuracy remains high thanks to sequence aggregation and segregated invariance-equivariance in our model.
+Conditioning only the transformer or only the predictor leads to intermediate results, with predictor conditioning proving more critical for equivariance.
+We also vary the dimensionality of the learnable action embeddings: performance saturates around the default size of 128, with smaller sizes (e.g., 16 or 64) already sufficient to capture the rotation structure.
+Non-learnable embeddings significantly reduce equivariance, underscoring the importance of learned representations in the action pathway.
+Table 4: Ablation results for action conditioning (3DIEBench). All models use $M_{tr}=3$, $M_{val}=5$ (results over three random seeds).
+Variant Top-1 Acc. (%) Rotation ($R^{2}$)
 Act. conditioning
 None 87.36 $\pm$ 0.7 0.29 $\pm$ 0.04
-Random noise 58.16 $\pm$ 3.2 0.20 $\pm$ 0.07
-Ablated predictor act cond. 87.17 $\pm$ 0.3 0.37 $\pm$ 0.06
-Ablated transformer act cond. 86.33 $\pm$ 0.1 0.53 $\pm$ 0.05
-Act. embedding dim
-Non-learnable act emb 85.68 $\pm$ 0.2 0.34 $\pm$ 0.04
-$adim=16$ 86.29$\pm$ 0.4 0.70 $\pm$ 0.01
-$adim=64$ 87.11$\pm$ 0.2 0.70 $\pm$ 0.02
-$adim=128$ (default) 87.41 $\pm$ 0.5 0.71 $\pm$ 0.02
-$adim=256$ 87.26$\pm$ 0.6 0.72 $\pm$ 0.00
-Figure 8: Effect of training and inference sequence length on seq-JEPA’s performance; left:: Equivariant performance ($R^{2}$) for different training sequence lengths (inference sequence length fixed); middle: Classification performance on top of aggregate representations versus training sequence length (inference sequence length fixed); right: Classification performance on top of aggregate representations for different inference sequence lengths (training sequence length fixed)
-5.7 Role of training and inference sequence lengths
-In order to study the effect of training and inference sequence lengths on our framework’s performance, we trained seq-JEPA with different training sequence lengths and measured both equivariance of encoder representations and linear probe accuracy using different inference sequence lengths to form the aggregate representation (Figure 8).
-A general positive trend can be observed in equivariance performance for different transformations when increasing the training sequence length (Figure 8, left panel), which can be the impact of seeing more action-conditioned representations during training.
-For linear classification (invariance-related performance; Figure 8, middle panel), the models trained on 3DIEBench and via predictive learning across saccades benefit from a longer training sequence length.
-However, in the augmentation setting with CIFAR100, a small drop in accuracy is observed for higher training sequence lengths.
-The observed differences across evaluation tasks likely stem from specific task requirements.
+No predictor cond. 87.17 $\pm$ 0.3 0.37 $\pm$ 0.06
+No transformer cond. 86.33 $\pm$ 0.1 0.53 $\pm$ 0.05
+Act. embedding dim.
+Non-learnable 85.68 $\pm$ 0.2 0.34 $\pm$ 0.04
+$a_{\text{dim}}=16$ 86.29 $\pm$ 0.4 0.70 $\pm$ 0.01
+$a_{\text{dim}}=64$ 87.11 $\pm$ 0.2 0.70 $\pm$ 0.02
+$a_{\text{dim}}=128$ (default) 87.41 $\pm$ 0.5 0.71 $\pm$ 0.02
+$a_{\text{dim}}=256$ 87.26 $\pm$ 0.6 0.72 $\pm$ 0.00
+Figure 7: Effect of training and inference sequence length on seq-JEPA’s performance; left:: Equivariant performance ($R^{2}$) versus training sequence length; middle: Classification accuracy versus training sequence length; right: Classification accuracy versus inference sequence length.
+4.6 Scaling Properties: Role of Training and Inference Sequence Lengths
+We study the effect of both training and inference sequence lengths on performance across tasks, i.e., scalability of seq-JEPA in terms of context length (Figure 7).
+We draw inspiration from recent findings in large-scale foundation models, where increased training and inference context—whether in text (Brown et al., 2020; Touvron et al., 2023), vision (Zellers et al., 2021; Chen et al., 2021), or video (Bain et al., 2021; Arnab et al., 2021)—consistently leads to stronger representations.
+We observe that equivariance generally improves with longer training sequences (left panel), likely due to increased exposure to action-conditioned transitions.
+Classification performance on 3DIEBench and STL-10 (middle panel) similarly benefits.
+In contrast, on CIFAR100 and Tiny ImageNet with synthetic augmentations, longer training sequences slightly decrease classification performance.
 We hypothesize that leveraging and aggregating a sequence of action-observation pairs, i.e.
-seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observation.
-In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased sequence length.
+seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observations.
+In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased training sequence length.
 Similarly, in the case of predictive learning across saccades, each eye movement and its subsequent glance provides additional information that can be leveraged for learning a richer aggregate representation.
-Such an advantage may not be present when using artificially-augmented images which can also explain the classification performance gap between seq-JEPA and the best-performing baselines in this setting (Table 2).
-Finally, during inference (Figure 8, right panel), all transformation settings benefit from more views and a richer context to improve the quality of aggregate representation.
+At inference time, all datasets benefit from longer context lengths ($M_{val}$), confirming that richer aggregate representations yield stronger performance (right plot).
+This scalability via sequence lengths opens avenues for efficient representation learning with small foveated patches in lieu of full-frame inputs, mirroring how foundation models scale with input tokens at test time.
+Together with our transfer learning results on ImageNet-1k (Appendix B.7), these findings suggest that seq-JEPA’s architectural inductive bias enables graceful scaling via longer sequence lengths.
+5 Related Work
+Non-Generative World Models and Joint-Embedding Predictive Architectures.
+Non-generative world models predict the effect of transformations or actions directly in latent space, avoiding reconstruction in pixel space.
+This includes contrastive SSL methods that model transformed views from context representations (van den Oord et al., 2019; Gupta et al., 2024), as well as approaches in model-based reinforcement learning (RL) to improve sample efficiency, generate intrinsic rewards, or capture environment transitions (Schwarzer et al., 2021; Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023; Guo et al., 2022).
+Joint-embedding predictive architectures (LeCun, 2022) form a structured subclass of non-generative world models.
+Rather than aligning or contrasting representations directly, they introduce an asymmetric predictor conditioned on transformation parameters to infer the outcome of an action applied to a latent view.
+Examples include I-JEPA(Assran et al., 2023), which predicts masked regions from positional cues, and IWM (Garrido et al., 2024), which conditions on augmentation parameters.
+JEPAs have been recently extended to physical reasoning (Garrido et al., 2025) and offline planning (Sobal et al., 2025), illustrating the framework’s versatility in structured predictive representation learning.
+Equivariant SSL.
+Equivariant SSL methods aim to retain transformation-specific information in the latent space, typically by augmenting invariant objectives with an additional equivariance term.
+Some approaches directly predict transformation parameters (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024; Dangovski et al., 2022).
+Methods such as EquiMod (Devillers and Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space via a predictor in addition to their invariant objective.
+SEN (Park et al., 2022) similarly predicts transformed representations but omits the invariance term.
+Xiao et al.
+(2021) use contrastive learning with separate projection heads for each augmentation, treating same-augmentation pairs as negatives.
+ContextSSL (Gupta et al., 2024) conditions representations on both current actions and recent context using a memory module and employs a dual predictor for transformation prediction to avoid collapsing to invariance.
+Other approaches (Shakerinava et al., 2022; Gupta et al., 2023; Yerxa et al., 2024) do not require explicit transformation parameters but instead enforce equivariance by applying the same transformation to multiple view pairs and minimizing a distance-based loss.
+Finally, action-conditioned JEPAs like IWM incorporate augmentation parameters and mask positions by conditioning the predictor to induce equivariance without additional objectives, though they still face the invariance-equivariance tradeoff depending on predictor capacity (Garrido et al., 2024).
+Positioning of Our Work.
+seq-JEPA belongs to the family of joint-embedding predictive architectures and is a non-generative world model.
+Unlike most equivariant SSL methods, seq-JEPA does not rely on an equivariance loss or transformation prediction objective, nor does it require view pairs with matched transformations.
+Instead, it introduces architectural inductive biases to JEPAs (Assran et al., 2023; Garrido et al., 2024) to learn separate invariant and equivariant representations, which emerge naturally from its action-conditioned sequential predictive learning.
+Furthermore, in contrast to ContextSSL, which extends the two-view contrastive setting of contrastive predictive coding van den Oord et al.
+(2019) using a transformer decoder projector conditioned on previous views via key-value caching, seq-JEPA operates on sequences of action-observation pairs in an online end-to-end manner by incorporating a sequence model (e.g.
+a transformer encoder) as a learned working memory during both training and inference.
+Moreover, while ContextSSL aims to adapt equivariance to recent transformations by dynamically modifying the training distribution, seq-JEPA is designed to explicitly learn both invariant and equivariant representations with respect to a specified set of transformations, and is also well-suited for downstream tasks that require multi-step observation aggregation (e.g., predictive learning across saccades and path integration).
 6 Limitations and Future Perspectives
-In this paper, we proposed seq-JEPA, a world modeling paradigm for learning architecturally segregated transformation-invariant and -equivariant representations.
-Although we have shown the feasibility of our world model for multiple transformation settings in image modality, we have not yet explored other modalities such as video, text, or multi-modal data.
-Given the architecture of seq-JEPA, it is possible to utilize the sequence aggregator to fuse representations from multiple modalities.
-Another promising direction is to build efficient world models for embodied agents with egocentric vision via scaling our approach to larger image and video datasets and take advantage of a sequence of foveated observations based on salience in a setup similar to predictive learning across saccades.
-Our transfer learning results on ImageNet-1k in Section B.6 hints at the possibility of building such a world model.
-Finally, similar to other equivariant SSL techniques, seq-JEPA requires knowledge of transformation groups.
-Developing a group-agnostic equivariant method is a challenge that remains largely unexplored in SSL.
-Acknowledgements
+We have validated viability of seq-JEPA across a range of transformations in the image domain.
+Here, we discuss limitations and possible future directions.
+First, the transformer-based aggregator in seq-JEPA could support multi-modal fusion across language, audio, or proprioceptive inputs—enabling multi-modal world modeling and generalization.
+Second, while our results show clear benefits from longer sequences, we have experimented with relatively short training and inference sequence lengths as the ResNet backbone is also trained end-to-end.
+The sequence scaling trends observed in the paper suggest that seq-JEPA can benefit from longer context windows over representations of pre-trained backbones as in (Pang et al., 2023; Lin et al., 2024).
+Third, our method assumes access to a known transformation group (e.g., $SO(3)$ for 3D rotations).
+Designing group-agnostic or learned transformation models (Finzi et al., 2021) without access to transformation parameters or pairs of same transformation is an open challenge in equivariant SSL that future work may tackle.
+Finally, our preliminary ImageNet-1k transfer results (Appendix B.7) point to potential for broader generalization.
+Scaling seq-JEPA to larger foveated image settings or video and multi-modal datasets such as Song et al.
+(2023) could support the development of lightweight, saliency-driven agents capable of learning efficiently from partial observations in embodied settings with a limited field of vision.
+Acknowledgments and Disclosure of Funding
 This project was supported by funding from NSERC (Discovery Grants RGPIN-2022-05033 to E.B.M., and RGPIN-2023-03875 to S.B.), Canada CIFAR AI Chairs Program and Google to E.B.M., Canada Excellence Research Chairs (CERC) Program, Mila - Quebec AI Institute, Institute for Data Valorization (IVADO), CHU Sainte-Justine Research Centre, Fonds de Recherche du Québec–Santé (FRQS), and a Canada Foundation for Innovation John R.
@@
 The original $256\times 256$ images are resized to a $128\times 128$ resolution for all experiments.
-Normalization is done using the means and standard deviations provided by Garrido et al.
-(2023), i.e., $\mu=[0.5016,0.5037,0.5060]$ and $\sigma=[0.1030,0.0999,0.0969]$ for the three RGB channels, respectively.
+Normalization is done using the means and standard deviations in Garrido et al.
+[2023], i.e., $\mu=[0.5016,0.5037,0.5060]$ and $\sigma=[0.1030,0.0999,0.0969]$ for the three RGB channels, respectively.
 CIFAR100.
@@
 STL10.
-In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE (Linardos et al., 2021), resize the output saliencies back to $96\times 96$, and store them alongside original images.
-We use normalization parameters $\mu=[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively.
-After sampling fixations from saliencies, the patches are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$).
+In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE [Linardos et al., 2021], resize the output saliencies back to $96\times 96$, and store them alongside original images.
+We use normalization parameters $\mu[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively.
+After sampling fixations from saliencies, the patches that are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$).
+For IoR, we zero-out a circular area with radius of $16$ around each previous fixation.
 Transformation parameters.
-For 3DIEBench, we use the rotation and color parameters provided with images for action conditioning as done in (Garrido et al., 2023).
+For 3DIEBench, we use the rotation and color parameters provided with images for action conditioning as done in Garrido et al.
+[2023].
 For the augmentation setting, we use the parameters corresponding to each of the three augmentations and form the action as the relative augmentation vector between two images.
@@
 A.2 Training and evaluation details
+Additional Training Details.
+We used the PyTorch framework for training all models.
+For experiments that use CIFAR100 and low-resolution STL-10 patches in predictive learning across saccades, we use the CIFAR variant of ResNet-18.
+For models trained with AdamW, we used with default $\beta_{1}$ and $\beta_{2}$, a weight decay of $0.001$, and a learning rate of $4\times 10^{-4}$ with a linear warmup for $20$ epochs starting from $10^{-5}$ followed by a cosine decay back to $10^{-5}$.
+For models trained with Adam, we use the Adam optimizer with a learning rate of $10^{-3}$, default $\beta_{1}$ and $\beta_{2}$, and no weight decay.
 Protocols for training evaluation heads.
 For linear probing, we follow a common SSL protocol and train a linear classifier on top of frozen representations with a batch size of 256 for 300 epochs using the Adam optimizer with default hyperparameters.
-For action prediction, we follow a similar protocol as SIE (Garrido et al., 2023).
+For action prediction, we follow a similar protocol as SIE [Garrido et al., 2023].
 Specifically, for rotation, color jitter, and crop, we train an MLP regressor with a hidden dimension of 1024 and ReLU activation for 300 epochs.
@@
 The predictor has a hidden dimension of 512-d with ReLU activation.
-We use the same EMA setup outlined in the original paper (Grill et al., 2020), i.e., the EMA parameter $\tau$ starts from $\tau_{\text{base}}=0.996$ and is increased following a cosine schedule.
+We use the same EMA setup outlined in the original paper [Guo et al., 2020], i.e., the EMA parameter $\tau$ starts from $\tau_{\text{base}}=0.996$ and is increased following a cosine schedule.
 SimCLR.
@@
 B.1 Evaluation results on 3DIEBench for models conditioned on rotation and color
-We further provide evaluation results for seq-JEPA and equivariant baselines on 3DIEBench conditioned on both rotation and color transformations in Table 5.
-The results show that all methods suffer a drop in classification performance and become highly sensitive to color.
-Similar performance degradations have been previously observed with 3DIEBench (Garrido et al., 2023; Gupta et al., 2024), yet without a convincing explanation.
-One possible explanation in line with Principle II in Wang et al.
-(2024) is that in 3DIEBench, floor and light hue are fairly decoupled from class information (low class relevance).
+Table 5 reports performance of seq-JEPA and several equivariant baselines when conditioned on both rotation and color in 3DIEBench.
+All methods suffer a drop in classification performance and become highly sensitive to color.
+Similar performance degradations have been previously observed with 3DIEBench [Garrido et al., 2023, Gupta et al., 2024], though without a clear explanation.
+One plausible explanation aligned with Principle II in Wang et al.
+[2024], is that color in 3DIEBench, i.e., floor and light hue, are weakly correlated with class labels in 3DIEBench (low class relevance).
 Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
-Table 5: Quantitative evaluation of learned representations on 3DIEBench for rotation and color prediction (equivariance) and linear probe classification (invariance).
-All models are conditioned on both rotation and color.
+Table 5: Evaluation on 3DIEBench for rotation and color prediction (equivariance) and linear probe classification (invariance). All models are conditioned on both rotation and color.
 Method Classification (top-1) Rotation pred. ($R^{2}$) Color pred. ($R^{2}$)
@@
 B.2 Measuring invariance of aggregate representations
-In order to have a sense of invariance of the aggregate representation, we tried to predict the individual rotation parameters of each view for a seq-JEPA model with $M_{val}=3$ trained on 3DIEBench and conditioned on rotation.
-To avoid leaking the rotation parameters to $z_{AGG}$ during prediction, we zero out the action embeddings.
-The mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; which indicates that the aggregate representation has not encoded rotation in any of the views.
+In order to have a more direct measure of invariance of the aggregate representation, we trained a model on 3DIEBench with $M_{val}=3$ and rotation conditioning, then attempted to predict the individual rotation parameters of each input view from $z_{\text{AGG}}$.
+To prevent leakage of rotation information during prediction, all action embeddings were zeroed out.
+The resulting mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; indicating that the aggregate representation contains little to no rotation-specific information.
 B.3 Complete evaluation results for linear probing on top of aggregate representations
-In this section, we provide our complete evaluation results for linear probing on top of seq-JEPA’s aggregate representations for three transformation sets with different training and inference sequence lengths.
+We provide our complete evaluation results for linear probing on top of seq-JEPA’s aggregate representations for our three transformation settings with different training and inference sequence lengths.
 Figure 8 shows the top-1 accuracy on 3DIEBench models conditioned on rotation.
 Figure 9 shows top-1 accuracy on STL-10 for models trained via predictive learning across saccades.
-Figures 11 and 12 show top-1 accuracy on CIFAR100 and Tiny ImageNet, respectively, with different types of action conditioning (crop, color jitter, blur, or all three).
-We can see the general trends of Figure 8 discussed in Section 5.7 also in these heatmaps.
+Figures 10 and 11 show top-1 accuracy on CIFAR100 and Tiny ImageNet, respectively, with different types of action conditioning (crop, color jitter, blur, or all three).
+These heatmaps reflect the same trends observed in Figure 7 and discussed in Section 4.6, illustrating the consistent effect of sequence length on representation quality.
 Figure 8: seq-JEPA’s performance on 3DIEBench with rotation conditioning; the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
@@
 For completeness, we provide linear probe classification on encoder representations for different transformation settings in Table 6 and compare them with accuracy on aggregate representations for different inference evaluation lengths.
-The aggregate representation generally achieves a higher classification performance thanks to the architectural inductive bias in seq-JEPA.
+The aggregate representation generally achieves a much higher classification performance thanks to the architectural inductive bias in seq-JEPA.
 Table 6: Comparison of seq-JEPA classification performance across datasets and conditioning.
@@
 3DIEBench Rotation 1 84.88 84.08 85.34 85.31
-3DIEBench Rotation 3 82.49 81.72 86.14 87.41
+3DIEBench Rotation 3 82.49 81.72 85.32 87.41
 3DIEBench Rotation + Color 4 74.88 71.14 75.97 79.31
@@
 Tiny ImageNet Blur 3 27.20 28.83 34.82 35.79
-STL-10 None 4 61.38 62.21 69.06 71.12
-STL-10 Position 4 81.20 71.45 81.53 83.67
-STL-10 Position (no saliency) 4 79.29 63.14 76.93 80.09
-STL-10 Position (no IoR) 4 72.49 68.95 76.84 78.21
-B.5 Visualization of UMAP projections for models without action conditioning
-In Figure 13 we provide the UMAP projections for seq-JEPA representations trained on 3DIEBench but without action conditioning.
-Figure 14 shows UMAP projections for seq-JEPA representations trained on STL-10 via predictive learning across saccades without action conditioning.
-In these visualizations, the color gradients that implied high equivariance in UMAPs of action-conditioned models are either absent or less pronounced.
-Figure 13: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations.
-The model is trained on 3DIEBench without any action conditioning with $M_{tr}=3$ and $M_{val}=5$.
-Encoder representations for each view observation, color-coded by class category (left) and rotation angle (middle).
-The aggregate token representation for $M_{val}=5$ observations during inference, color-coded by class category (right).
-Figure 14: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations.
-The model is trained on STL-10 via predictive learning across saccades with no action conditining and with $M_{tr,val}=4$.
-Encoder representations for each view observation, color-coded by class category (left) and $X$ and $Y$ pixel coordinates (middle panels).
-The aggregate token representation for $M_{val}=4$ observation patches during inference, color-coded by class category (right).
-B.6 Transfer learning results on ImageNet-1k
-In this section, we consider transfer learning on ImageNet-1k using our frozen model trained on STL-10 via predictive learning across saccades.
-We follow the same evaluation protocol as the in-distribution setting and train a linear probe on ImageNet-1k training set and evaluate it on the validation set using the aggregate representation for different inference sequence lengths.
-We use DeepGaze-IIE to extract saliency maps of ImageNet-1k, resize them to $224\times 224$, and use them to extract the foveated patches.
-Figure 15 shows the top-1 accuracy on ImageNet-1k for two different patch sizes (32 and 84).
-These results hint at the possibility of scaling seq-JEPA to larger data regimes by striking a trade-off between the patch size and sequence length.
-For example, using full-resolution $224\times 224$ images, a forward pass through a ResNet-18 encoder takes approximately $1.83\times 10^{9}$ flops, while a forward pass of seq-JEPA with the default transformer encoder, $M_{val}=4$ and $84\times 84$ patches takes $1.25\times 10^{9}$ flops.
-Using the CIFAR variant of ResNet-18 with $32\times 32$ patches and $M_{val}=8$ requires $2.35\times 10^{9}$ flops.
-Figure 15: Linear probe transfer learning accuracy on ImageNet-1k for two different patch sizes; the model is trained on STL-10 via predictive learning across saccades.
+STL-10 None 4 61.38 62.21 69.06 70.45
+STL-10 Position 4 81.20 71.45 81.53 83.44
+STL-10 Position (no saliency) 4 79.29 63.14 76.93 79.85
+STL-10 Position (no IoR) 4 72.49 68.95 76.84 77.97
+B.5 Additional UMAP Visualizations
+In Figure 12, we visualize the UMAP projections of seq-JEPA representations trained on 3DIEBench without action conditioning.
+Similarly, Figure 14 shows projections for models trained on STL-10 via predictive learning across saccades, also without action conditioning.
+Compared to the action-conditioned counterparts (Figures 5 and 13), these projections exhibit weaker or no smooth color gradients in their corresponding transformation-colored UMAP—indicating reduced equivariance to transformation parameters.
+Figure 12: 2-D UMAP projections of seq-JEPA representations on 3DIEBench without action conditioning ($M_{tr}=3$, $M_{val}=5$).
+Encoder outputs are color-coded by class (left) and rotation angle (middle); aggregate token representations are color-coded by class (right).
+Figure 13: 2-D UMAP projections of seq-JEPA representations on STL-10 with action conditioning ($M_{tr}=M_{val}=4$).
+Encoder outputs are color-coded by class (top-left) and by $X$/$Y$ fixation coordinates (bottom); the aggregate token is color-coded by class (top-right).
+Figure 14: 2-D UMAP projections of seq-JEPA representations on STL-10 without action conditioning ($M_{tr}=M_{val}=4$).
+Encoder outputs are color-coded by class (left) and fixation coordinates (middle); the aggregate token is color-coded by class (right).
+B.6 Details of Path Integration Experiments.
+While an agent executes a sequence of actions in an environment, transitioning from an initial state to a final state, it should be capable of tracking its position by integrating its own actions.
+This is also a crucial cognitive ability that enables animals to estimate their current state in their habitat [McNaughton et al., 2006].
+Here, we evaluate whether seq-JEPA is capable of path integration.
+Given the sequence of observations $\{x_{i}\}_{i=1}^{M+1}$ generated from transformations $\{t_{i}\}_{i=1}^{M+1}$ and the corresponding relative action embeddings $\{a_{i}\}_{i=1}^{M}$, we define the task of path integration over the sequence of actions as predicting the relative action that would directly transform $x_{1}$ to $x_{M+1}$ given $z_{AGG}$ and $a_{M}$.
+In other words, given the aggregate representation of a sequence of action-observation pairs and the next action, we would like to predict the overall position change from the starting point ($x_{1}$) to the end point ($x_{M+1}$).
+We consider path integration for rotation angles in 3DIEBench and across eye movements with STL-10.
+For rotations, the task is integrating a series of object rotations from the first view to the last, i.e.
+angular path integration.
+For eye movements, the task is integrating the eye movements from the first saccade to the last, i.e.
+visual path integration.
+To measure path integration performance for inference sequence length $M$, we train a regression head on top of the concatenation of $z_{AGG}$ and $a_{M}$ to predict the transformation from $x_{1}$ to $x_{M+1}$.
+Figure 6 shows that seq-JEPA performs well in both angular and visual path integration.
+The red curve corresponds to the performance of the original seq-JEPA.
+The blue curve corresponds to experiments in which the action embeddings are ablated (zeroed-out during inference for all views).
+The green curve corresponds to experiments in which the encoder (visual) representations are ablated during inference.
+As expected, path integration becomes more difficult as the number of observations increases (red curves).
+Ablating action conditioning (blue curves) results in failure of path integration.
+On the other hand, ablating the visual representations (green curves) results only in a small performance drop compared to the original model, indicating that action conditioning is the key factor that enables path integration.
+B.7 Transfer learning results on ImageNet-1k
+To evaluate generalization of our model beyond STL-10, we assess transfer performance of the model trained on STL-10 patches via predictive learning across saccades on ImageNet-1k.
+We follow the same linear probing protocol as in-distribution evaluations: we freeze the ResNet and transformer encoder and train a linear classifier on aggregate representations generated from foveated patches.
+We extract saliency maps for ImageNet-1k images using DeepGaze IIE, resize them to $224\times 224$, and sample sequences of patches sized $32\times 32$ or $84\times 84$ for this out-of-distribution (OOD) evaluation setting.
+Figure 15 shows top-1 linear probe accuracy on ImageNet-1k validation set across varying inference sequence lengths ($M_{val}$).
+For both patch sizes, performance improves with longer inference sequences, validating seq-JEPA’s ability to benefit from extended context even in this difficult OOD ImageNet-1k setting.
+These results echo findings in the main experiments for different training and inference sequence lengths and confirming the possibility of model’s scalability in terms of data, parameter count, and compute.
+Figure 15: Linear probe transfer learning accuracy on ImageNet-1k for two different patch sizes; the model is pre-trained on STL-10 via predictive learning across saccades.

### v3

@@
 Abstract
-Current self-supervised algorithms commonly rely on transformations such as data augmentation and masking to learn visual representations.
-This is achieved by enforcing invariance or equivariance with respect to these transformations after encoding two views of an image.
-This dominant two-view paradigm often limits the flexibility of learned representations for downstream adaptation by creating performance trade-offs between high-level invariance-demanding tasks such as image classification and more fine-grained equivariance-related tasks.
-In this work, we proposes seq-JEPA, a world modeling framework that introduces architectural inductive biases into joint-embedding predictive architectures to resolve this trade-off.
-Without relying on dual equivariance predictors or loss terms, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to specified transformations and another invariant to them.
+Joint-embedding self-supervised learning (SSL) commonly relies on transformations such as data augmentation and masking to learn visual representations, a task achieved by enforcing invariance or equivariance with respect to these transformations applied to two views of an image.
+This dominant two-view paradigm in SSL often limits the flexibility of learned representations for downstream adaptation by creating performance trade-offs between high-level invariance-demanding tasks such as image classification and more fine-grained equivariance-related tasks.
+In this work, we propose seq-JEPA, a world modeling framework that introduces architectural inductive biases into joint-embedding predictive architectures to resolve this trade-off.
+Without relying on dual equivariance predictors or loss terms, seq-JEPA simultaneously learns two architecturally separate representations for equivariance- and invariance-demanding tasks.
 To do so, our model processes short sequences of different views (observations) of inputs.
@@
 A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation.
-Empirically, seq-JEPA demonstrates strong performance on both equivariant and invariant benchmarks without sacrificing one for the other.
+Empirically, seq-JEPA demonstrates strong performance on both equivariance- and invariance-demanding downstream tasks without sacrificing one for the other.
 Furthermore, it excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
@@
 While some equivariant SSL approaches have reported minor gains on tasks typically associated with invariance (e.g., classification) (Devillers and Lefort, 2022; Park et al., 2022; Gupta et al., 2023), a growing body of work highlights a fundamental trade-off between learning invariance and equivariance, i.e., models that capture equivariance-related style latents do not fare well in classification and vice versa (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024; Rusak et al., 2025).
-This trade-off has recently received theoretical support (Wang et al., 2024), underscoring the need for new architectural and objective designs that can reconcile these competing goals.
+This trade-off has recently received theoretical support (Wang et al., 2024b), underscoring the need for new architectural and objective designs that can reconcile these competing goals.
 In contrast to the two-view paradigm in SSL, humans and other animals rely on a sequence of actions and consequent observations (views) for developing appropriate visual representation during novel object learning  (Harman et al., 1999; Vuilleumier et al., 2002).
@@
 Through ablations, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network.
-In contrast, the aggregate representation of views, produced at the output of the transformer, becomes largely transformation/action-invariant.
-This emergent architectural disentanglement of invariant and equivariant representations is central to seq-JEPA’s competitive performance compared to invariant and equivariant SSL methods on both categories of tasks (Figure 3).
-Unlike most prior equivariant SSL methods (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), our model does not rely on explicitly crafted loss terms or objectives to achieve equivariance, nor is it instructed to learn the decomposition of two representation types.
+In contrast, the aggregate representation of views, produced at the output of the transformer proves highly effective for invariance-demanding downstream tasks.
+This emergent architectural disentanglement of invariance and equivariance is central to seq-JEPA’s competitive performance compared to invariant and equivariant SSL methods on both categories of tasks (Figure 3).
+Unlike most prior equivariant SSL methods (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), our model does not rely on explicitly crafted loss terms or objectives to achieve equivariance, nor is it instructed to learn the decomposition of two representations.
 Instead, the dual representation structure arises naturally from the model architecture and action-conditioned predictive learning.
-Beyond resolving the invariance-equivariance trade-off, seq-JEPA further benefits from moving away from the two-view paradigm in SSL.
-We show that it performs well on tasks requiring integration over sequences of observations.
+Beyond resolving the invariance-equivariance trade-off, seq-JEPA further benefits from processing a sequence of observation views; we show that it performs well on tasks requiring integration over sequences of observations.
 In one scenario, inspired by embodied vision in primates, our model learns image representations without augmentations or masking, solely by predicting across simulated eye movements (saccades).
@@
 Our key contributions are as follows:
-  • We introduce seq-JEPA, a self-supervised world model that learns architecturally distinct equivariant and invariant representations through sequential prediction over action-observation pairs, without requiring explicit equivariance losses or dual predictors.
+  • We introduce seq-JEPA, a self-supervised world model that learns architecturally distinct representations for invariance- and equivariance-demanding downstream tasks through sequential prediction over action-observation pairs, without requiring explicit equivariance losses or dual predictors.
   • We empirically validate that seq-JEPA matches or outperforms existing invariant and equivariant SSL methods across tasks requiring either representational property.
@@
 Except for $z_{M}$, each $z_{i}$ is concatenated with its corresponding action embedding and passed to a transformer encoder $g$ (no MLP projector is used after the encoder), which aggregates the sequence of action-observation pairs.
-The transformer uses a learnable [AGG] token (analogous to the [CLS] token in ViTs (Dosovitskiy et al., 2020)) to generate the aggregate representation:
+The transformer uses a learnable [AGG] token (analogous to the [CLS] token in ViT (Dosovitskiy et al., 2020)) to generate the aggregate representation:
 $z_{AGG}=g((z_{1},a_{1}),(z_{2},a_{2})...,(z_{M-1},a_{M-1}),z_{M})$ (3)
@@
 The ground truth $z_{M+1}$ is computed using a target encoder—an exponential moving average (EMA) of $f$.
-The target representation is passed through a stop-gradient operator (sg) to avoid representational collapse.
-The training objective is to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$:
+This target representation is passed through a stop-gradient operator (sg) to avoid representational collapse (Grill et al., 2020; Chen and He, 2021; Assran et al., 2023).
+The training objective is to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$ with the loss function:
 $\mathcal{L}_{\mathrm{seq-JEPA}}=1-\frac{\hat{z}_{M+1}}{\left\lVert{\hat{z}_{M+1}}\right\rVert_{2}}\cdot\frac{\texttt{sg}(z_{M+1})}{\left\lVert{\texttt{sg}(z_{M+1})}\right\rVert_{2}}.$ (5)
 No additional loss terms or equivariance-specific predictors are used during training.
-Figure 1: seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn architecturally distinct invariant and equivariant representations.
+Figure 1: seq-JEPA is a self-supervised world model that leverages a sequence of action-observation pairs to learn architecturally distinct representations for downstream tasks requiring transformation invariance or equivariance.
 Figure 2: a.
@@
 2.3 Action and observation sets
-To evaluate generalization across transformation types, we define three sets of action-observation pairs (Figure 2). See Appendix A.1 for details of each setup.
+To evaluate generalization across transformation types, we consider three sets of action-observation pairs (Figure 2). See Appendix A.1 for details of each setup.
 3D Invariant Equivariant Benchmark (3DIEBench).
-We first evaluate on the 3DIEBench dataset (Garrido et al., 2023), designed for equivariance analysis.
+The 3DIEBench dataset (Garrido et al., 2023) is designed to evaluate representational invariance and equivariance.
 It includes 3D object renderings with variations in rotation, floor hue, and lighting.
+In this benchmark, the action between two views is the relative difference of these three factors of variation.
 We primarily study equivariance to $SO(3)$ rotations and secondarily to appearance factors of floor and light hue.
 Hand-Crafted Augmentations.
-In this setting, we use transformed views generated via common SSL augmentations (e.g., crop, color jitter, blur), where actions correspond to relative augmentation parameters.
-We use CIFAR100 and Tiny ImageNet, and follow EquiMod’s augmentation protocol (Devillers and Lefort, 2022).
+In this setting, we use transformed views generated via common SSL augmentations (e.g., crop, color jitter, blur), and actions correspond to relative augmentation parameters.
+We use CIFAR100 and Tiny ImageNet for experiments in this setup, and follow EquiMod’s augmentation protocol (Devillers and Lefort, 2022).
 Predictive Learning Across Saccades (PLS).
@@
 Our PLS has a similar flavor to I-JEPA (Assran et al., 2023) but does not require engineered masking strategies.
-In PLS, we train seq-JEPA on sequences of patches extracted from high-resolution images.
-For instance, with STL-10, we use $32\times 32$ patches to form the observation sequence.
+In PLS, we train seq-JEPA on sequences of patches extracted from full-resolution images.
+For instance, with STL-10 dataset, we use small $32\times 32$ patches to form the observation sequence.
 In this setting, actions correspond to the relative positions between patch centers, simulating saccadic eye movements and inducing 2-D positional equivariance to representations.
@@
 We compare seq-JEPA against both invariant and equivariant SSL baselines.
-Invariant methods include SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022).
+Invariant methods include SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022) and VICReg with trajectory regularization (Wang et al., 2024a).
 Equivariant methods include SEN (Park et al., 2022), EquiMod (Devillers and Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024).
@@
 The predictor is a 2-layer MLP with 1024 hidden units and ReLU activation.
-In order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train baselines that typically use MLP projectors in two variants: (1) with original MLP projector; and (2) with the MLP replaced by a transformer encoder and a sequence length of one.
-We report the better-performing variant to ensure a fair comparison.
+In order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we trained baselines that typically use MLP projectors in two variants: (1) with original MLP projector; and (2) with the MLP replaced by a transformer encoder and a sequence length of one.
+We did not see any benefit from switching to transformer projectors in any of the baselines, and include the transformer-projector results in Appendix B.2.
 All models are trained from scratch with a batch size of 512.
@@
 We report the $R^{2}$ score on the test set.
+In addition to action decoding $R^{2}$, we also report retrieval-based metrics including Mean Reciprocal Rank (MRR), Hit@1, and Hit@5 (Kipf et al., 2020; Park et al., 2022; Garrido et al., 2023) to evaluate the quality of the predictor.
 As a proxy measure of invariance, we use top-1 classification accuracy of a linear probe on top of frozen representations.
@@
 4.1 Quantitative evaluation on 3DIEBench
-We use the 3DIEBench benchmark to quantitatively compare invariant and equivariant representations in seq-JEPA with baseline methods.
+We use the 3DIEBench benchmark to quantitatively compare performance on equivariance- and invariance-demanding downstream tasks in seq-JEPA with baseline methods.
 This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object classification.
-Table 3 provides a summary of our evaluation on the 3DIEBench where equivariant methods have been conditioned on rotation.
+Table 3 provides a summary of our evaluation on the 3DIEBench where equivariant methods have been conditioned on rotation (a 4-D quaternion representing the relative rotation between two views).
 In addition to the relative rotation between two views, in the last column we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view.
@@
 Among invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance.
+Adding a linear trajectory regularization loss to VICReg Wang et al.
+(2024a) without using ground-truth transformations improves over VICReg, which shows that imposing geometric priors can improve both invariant and equivariant performance—even when these priors do not fully materialize in the environment, e.g., when we have non-smooth angle changes as in 3DIEBench.
 Among equivariant baselines, SIE and ContextSSL yield strong rotation prediction performance due to their specialized equivariance predictors and loss functions, but underperform in classification.
@@
 In contrast, seq-JEPA achieves strong performance in both, matching the best rotation $R^{2}$ scores while exceeding baselines in classification, with gains increasing with inference sequence length $M_{val}$.
-Ablating action conditioning leads to a sharp drop in equivariance but retains classification accuracy, confirming our hypothesis that action-conditioned sequential aggregation enables a segregated invariant-equivariant representation structure.
-For additional results on 3DIEBench including models with varying training and inference sequence lengths and models conditioned on both rotation and color, see Appendices B.1 to B.3.
+Ablating action conditioning leads to a sharp drop in equivariance but retains classification accuracy, confirming our hypothesis that action-conditioned sequential aggregation enables a distinct representational structure for invariance- and equivariance-demanind tasks.
+For additional results on 3DIEBench including MRR, Hit@1, and Hit@5 metrics, performance of models with varying training and inference sequence lengths and models conditioned on both rotation and color, and evaluation on an out-of-distribution (OOD) set of 3DIEBench see Appendices B.1 to B.6.
 Figure 3: Top-1 linear classification (invariance) vs.
@@
 Subscripts indicate training and inference sequence lengths.
-Table 1: Evaluation on 3DIEBench for linear probe classification (invariance) and rotation prediction (equivariance).
+Table 1: Evaluation on 3DIEBench for linear probe classification (invariance-demanding) and rotation prediction (equivariance-demanding).
 Equivariant models and seq-JEPA are conditioned on rotation.
@@
 VICReg 80.48 0.20 0.36
+VICRegTraj 81.26 0.27 0.43
 Equivariant
@@
 4.3 Evaluation with Hand-Crafted Augmentations
-We assess invariance and equivariance under hand-crafted augmentations by training on CIFAR100 and Tiny ImageNet (Table 2).
+We assess downstream performance under hand-crafted augmentations by training on CIFAR100 and Tiny ImageNet (Table 2).
 Models with action conditioning are trained conditioned on crop, color jitter, blur, or all three (indicated in the first column of the table).
@@
 VICReg 61.35 0.49 0.11 -0.06 35.29 0.31 0.05 0.19
+VICRegTraj 61.07 0.46 0.14 0.02 34.95 0.28 0.09 0.16
 Equivariant
@@
 4.4 Predictive Learning across Saccades and Path Integration
-In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations to learn visual representations in a predictive manner.
-Table 6 shows that seq-JEPA reaches 83.44% top-1 accuracy on STL-10, comparable to SimCLR (85.23%) trained with full-resolution images and strong augmentations.
+In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations to learn visual representations.
+In Table 6, seq-JEPA reaches 83.44% top-1 accuracy on STL-10, comparable to SimCLR (85.23%) trained with full-resolution images and strong augmentations.
 This gap narrows further when increasing the inference sequence length from $M_{val}=4$ to $M_{val}=6$.
 Ablating action conditioning causes the accuracy on top of the aggregate representations to drop sharply, indicating that 2-D positional awareness is essential to forming semantic representations across simulated eye movements.
-Compared to Conv-JEPA, which accumulates prediction losses pairwise, seq-JEPA performs better in classification, highlighting the importance of sequence aggregation when dealing with partial observations in SSL.
+Compared to Conv-JEPA—which accumulates prediction losses pairwise—seq-JEPA performs better in classification, highlighting the importance of sequence aggregation when dealing with partial observations in SSL.
 Further ablations show that saliency-driven sampling and IoR are critical for forming informative, non-overlapping patch sequences and subsequently a high-quality aggregate representation.
 Interestingly, while random uniform patch sampling negatively impacts classification accuracy due to lower semantic content, it results in the highest positional equivariance as the model samples patches and corresponding saccade actions from a more diverse set of positions across the entire image, not just the salient regions.
-UMAP projections for PLS with and without action conditioning in Appendix B.5 further underscore the role of action conditioning in enabling positional equivariance.
+The UMAP projections for PLS with and without action conditioning in Appendix B.8 further underscore the role of action conditioning in enabling positional equivariance.
 Path Integration.
-In the context of eye movement-driven observations and in general sequential observations, another ability that naturally arises is path integration (McNaughton et al., 2006)—predicting the cumulative transformation/action from a sequence of actions.
+In the context of eye movement-driven or any sequential observations, an ability that naturally arises from predictive learning is path integration (McNaughton et al., 2006), i.e., predicting the cumulative transformation/action from a sequence of actions.
 We evaluate this task in both eye movements in PLS (visual path integration) and object rotations in 3DIEBench (angular path integration).
@@
 Figure 6: a) Visual path integration across eye movements, b) Angular path integration across object rotations (results over three random seeds)
-4.5 Ablations: Action Conditioning Is Key
+4.5 Action Conditioning Ablations
 To better understand the mechanisms underlying seq-JEPA’s invariant-equivariant representation learning and role of action conditioning, we perform a set of ablation experiments on 3DIEBench.
-Specifically, we study: (i) the role of action conditioning in the transformer and predictor; (ii) the impact of learning action embeddings and their dimensionality.
+Specifically, we study: (i) the role of action conditioning in the transformer and predictor; (ii) the impact of action embedding dimensionality.
 Table 7 summarizes our ablation results.
@@
 We also vary the dimensionality of the learnable action embeddings: performance saturates around the default size of 128, with smaller sizes (e.g., 16 or 64) already sufficient to capture the rotation structure.
-Non-learnable embeddings significantly reduce equivariance, underscoring the importance of learned representations in the action pathway.
 Table 4: Ablation results for action conditioning (3DIEBench). All models use $M_{tr}=3$, $M_{val}=5$ (results over three random seeds).
@@
 Act. embedding dim.
-Non-learnable 85.68 $\pm$ 0.2 0.34 $\pm$ 0.04
 $a_{\text{dim}}=16$ 86.29 $\pm$ 0.4 0.70 $\pm$ 0.01
@@
 We study the effect of both training and inference sequence lengths on performance across tasks, i.e., scalability of seq-JEPA in terms of context length (Figure 7).
-We draw inspiration from recent findings in large-scale foundation models, where increased training and inference context—whether in text (Brown et al., 2020; Touvron et al., 2023), vision (Zellers et al., 2021; Chen et al., 2021), or video (Bain et al., 2021; Arnab et al., 2021)—consistently leads to stronger representations.
-We observe that equivariance generally improves with longer training sequences (left panel), likely due to increased exposure to action-conditioned transitions.
-Classification performance on 3DIEBench and STL-10 (middle panel) similarly benefits.
+We draw inspiration from recent findings in foundation models, where increased training and inference context—whether in text (Brown et al., 2020; Touvron et al., 2023), vision (Zellers et al., 2021; Chen et al., 2021), or video (Bain et al., 2021; Arnab et al., 2021)—consistently leads to stronger representations.
+We observe that equivariance generally improves with longer training sequences (left panel).
+A possible explanation for this observation is the presence of more transitions $(z_{i},a_{i},z_{i+1})$ in working memory, which means the predictor has access to a richer context $z_{\mathrm{agg}}$.
+This enables the predictor to more accurately approximate the transition $p(z_{i+1}\mid z_{\mathrm{agg,i}},a_{i})$.
+Because accurate prediction requires the model to preserve and utilize information about $a_{t}$, the encoder is implicitly encouraged to learn structured, more equivariant representations.
+Classification performance on 3DIEBench and STL-10 (middle panel) benefits from a longer training sequence.
 In contrast, on CIFAR100 and Tiny ImageNet with synthetic augmentations, longer training sequences slightly decrease classification performance.
@@
 This includes contrastive SSL methods that model transformed views from context representations (van den Oord et al., 2019; Gupta et al., 2024), as well as approaches in model-based reinforcement learning (RL) to improve sample efficiency, generate intrinsic rewards, or capture environment transitions (Schwarzer et al., 2021; Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023; Guo et al., 2022).
-Joint-embedding predictive architectures (LeCun, 2022) form a structured subclass of non-generative world models.
-Rather than aligning or contrasting representations directly, they introduce an asymmetric predictor conditioned on transformation parameters to infer the outcome of an action applied to a latent view.
+Joint-embedding predictive architectures (LeCun, 2022) form a subclass of non-generative world models.
+They introduce an asymmetric predictor conditioned on transformation parameters to infer the outcome of an action applied to a latent view.
 Examples include I-JEPA(Assran et al., 2023), which predicts masked regions from positional cues, and IWM (Garrido et al., 2024), which conditions on augmentation parameters.
-JEPAs have been recently extended to physical reasoning (Garrido et al., 2025) and offline planning (Sobal et al., 2025), illustrating the framework’s versatility in structured predictive representation learning.
+JEPAs have been recently extended to physical reasoning (Garrido et al., 2025) and offline planning (Sobal et al., 2025), illustrating the framework’s versatility in representation learning and world modeling.
 Equivariant SSL.
@@
 (2021) use contrastive learning with separate projection heads for each augmentation, treating same-augmentation pairs as negatives.
-ContextSSL (Gupta et al., 2024) conditions representations on both current actions and recent context using a memory module and employs a dual predictor for transformation prediction to avoid collapsing to invariance.
+ContextSSL (Gupta et al., 2024) conditions representations on both current actions and recent context and employs a dual predictor for transformation prediction to avoid collapsing to invariance.
 Other approaches (Shakerinava et al., 2022; Gupta et al., 2023; Yerxa et al., 2024) do not require explicit transformation parameters but instead enforce equivariance by applying the same transformation to multiple view pairs and minimizing a distance-based loss.
-Finally, action-conditioned JEPAs like IWM incorporate augmentation parameters and mask positions by conditioning the predictor to induce equivariance without additional objectives, though they still face the invariance-equivariance tradeoff depending on predictor capacity (Garrido et al., 2024).
+Action-conditioned JEPAs incorporate augmentation parameters and mask positions by conditioning the predictor to induce equivariance without additional objectives (Garrido et al., 2024).
+Chavhan et al.
+(2023) use an ensemble of heads trained on top of a pre-trained SSL encoder to span a diverse spectrum of transformation sensitivities across the latent spaces of each head.
+Downstream probes then learn to linearly combine these feature heads depending on the desired invariance-equivariance trade-off.
 Positioning of Our Work.
@@
 Unlike most equivariant SSL methods, seq-JEPA does not rely on an equivariance loss or transformation prediction objective, nor does it require view pairs with matched transformations.
-Instead, it introduces architectural inductive biases to JEPAs (Assran et al., 2023; Garrido et al., 2024) to learn separate invariant and equivariant representations, which emerge naturally from its action-conditioned sequential predictive learning.
-Furthermore, in contrast to ContextSSL, which extends the two-view contrastive setting of contrastive predictive coding van den Oord et al.
+Instead, it leverages action conditioning and architectural inductive biases to learn two separate invariant and equivariant representations.
+In contrast to ContextSSL, which extends the two-view contrastive setting of contrastive predictive coding van den Oord et al.
 (2019) using a transformer decoder projector conditioned on previous views via key-value caching, seq-JEPA operates on sequences of action-observation pairs in an online end-to-end manner by incorporating a sequence model (e.g.
 a transformer encoder) as a learned working memory during both training and inference.
-Moreover, while ContextSSL aims to adapt equivariance to recent transformations by dynamically modifying the training distribution, seq-JEPA is designed to explicitly learn both invariant and equivariant representations with respect to a specified set of transformations, and is also well-suited for downstream tasks that require multi-step observation aggregation (e.g., predictive learning across saccades and path integration).
+Moreover, while ContextSSL aims to adapt equivariance to recent transformations by dynamically modifying the training distribution, seq-JEPA is designed to explicitly learn both invariant and equivariant representations with respect to a specified set of transformations, and is also well-suited for downstream tasks that require multi-step observation aggregation.
 6 Limitations and Future Perspectives
@@
 Here, we discuss limitations and possible future directions.
-First, the transformer-based aggregator in seq-JEPA could support multi-modal fusion across language, audio, or proprioceptive inputs—enabling multi-modal world modeling and generalization.
-Second, while our results show clear benefits from longer sequences, we have experimented with relatively short training and inference sequence lengths as the ResNet backbone is also trained end-to-end.
+First, seq-JEPA is capable of autoregressive prediction in time, and therefore, can be leveraged for autoregressive latent planning in control tasks.
+Second, the transformer-based aggregator in seq-JEPA could support multi-modal fusion across language, audio, or proprioceptive inputs—enabling multi-modal world modeling and generalization.
+Third, while our results show clear benefits from longer sequences, we have experimented with relatively short training and inference sequence lengths as the backbone is also trained end-to-end.
 The sequence scaling trends observed in the paper suggest that seq-JEPA can benefit from longer context windows over representations of pre-trained backbones as in (Pang et al., 2023; Lin et al., 2024).
-Third, our method assumes access to a known transformation group (e.g., $SO(3)$ for 3D rotations).
+Fourth, our method assumes access to a known transformation group (e.g., $SO(3)$ for 3D rotations).
 Designing group-agnostic or learned transformation models (Finzi et al., 2021) without access to transformation parameters or pairs of same transformation is an open challenge in equivariant SSL that future work may tackle.
 Finally, our preliminary ImageNet-1k transfer results (Appendix B.10) point to potential for broader generalization.
-Scaling seq-JEPA to larger foveated image settings or video and multi-modal datasets such as Song et al.
+Scaling seq-JEPA to larger foveated image settings or video and multi-modal datasets such as Ego4D Song et al.
 (2023) could support the development of lightweight, saliency-driven agents capable of learning efficiently from partial observations in embodied settings with a limited field of vision.
@@
 In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE [Linardos et al., 2021], resize the output saliencies back to $96\times 96$, and store them alongside original images.
-We use normalization parameters $\mu[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively.
+We use normalization parameters $\mu=[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively.
 After sampling fixations from saliencies, the patches that are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$).
@@
 All regression heads are trained using the Adam optimizer with default hyperparameters.
-Hardware. Each experiment was run on a single NVIDIA A100 GPU with 40GB of accelerator RAM. A single run of seq-JEPA on 3DIEBench with a training sequence length of three takes around 15 hours.
+Hardware. Each experiment was run on a single NVIDIA A100 GPU with 40GB of accelerator RAM.
+Compute cost and FLOP analysis.
+We report pretraining compute time across methods (see Table 5).
+All experiments are run under the same configuration: a single A100 GPU, $\times128$ resolution, batch size 512, for 1000 epochs on 3DIEBench.
+seq-JEPA with a sequence length of one has similar runtime to other baselines (e.g., BYOL, SimCLR), while seq-JEPA with sequence length of three incurs a moderate increase in wall-clock time (15.1 GPU-hours) due to encoding additional action-view pairs.
+Table 5: GPU-Hour comparison across methods
+Method A100 pretraining GPU hours
+SimCLR 11.4
+BYOL 11.1
+SIE 12.0
+VICReg 10.9
+EquiMod 11.8
+seq-JEPA (train seq len = 1) 12.3
+seq-JEPA (train seq len = 3) 15.1
+seq-JEPA’s inference cost is primarily governed by two factors: sequence length and input resolution.
+Below, we report detailed FLOP counts (in Gigaflops) per forward pass of a single datapoint across a range of configurations (Note: these reflect inference-time cost).
+Table 6: FLOPs for different seq-JEPA configurations (in gigaflops)
+Resolution / Config Encoder only Post-aggregator
+seq len = 1 seq len = 2 seq len = 3 seq len = 4 seq len = 8
+Each view is $\times128$ 0.60G 0.63G 1.24G 1.85G 2.46G 4.9G
+Views of $\times64$ vs. full image ($\times224$) 1.82G (224) / 0.15G (64) 0.18G 0.34G 0.51G 0.67G 1.32G
+Views of $\times84$, vs. full image ($\times224$) 1.82G (224) / 0.30G (84) 0.33G 0.64G 0.95G 1.26G 2.51G
+We observe that for single-view inference at $\times128$ resolution, the encoder requires only 0.60G FLOPs, which matches the compute cost of standard SSL baselines such as SimCLR and BYOL.
+As sequence length increases, the additional compute cost grows sub-linearly as additional views should be encoded and aggregated.
+For example, lengths 2 and 3—where invariant performance improves substantially—incur $\sim$1.2G and $\sim$1.85G FLOPs, respectively.
+The cost can be offset by reducing input resolution.
+In our saccade-based setup, we can feed low-resolution foveal glimpses (e.g., $\times64$ or $\times84$ crops sampled via saliency maps), which require only 0.15G–0.30G FLOPs per view.
+This enables the use of longer sequences at an overall compute cost comparable to full-resolution, single-view pipelines (e.g., compare the FLOPs of aggregating eight $\times64$ crops with encoding a single $\times224$ image in the second row of Table 6).
 A.3 Architectural details
@@
 We use $\lambda_{\text{inv}}=\lambda_{V}=10$, $\lambda_{C}=1$, and a projection head with 2048-2048-2048 intermediate dimensions.
+VICRegTraj.
+We use the same architecture as VICReg and add a trajectory loss (Equation 2 in Wang et al. [2024a]) with a coefficient $\lambda_{traj}=0.01$.
 Conditional BYOL.
@@
 One plausible explanation aligned with Principle II in Wang et al.
-[2024], is that color in 3DIEBench, i.e., floor and light hue, are weakly correlated with class labels in 3DIEBench (low class relevance).
+[2024b], is that color in 3DIEBench, i.e., floor and light hue, are weakly correlated with class labels in 3DIEBench (low class relevance).
 Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
@@
 seq-JEPA ($M_{tr}=4$, $M_{val}=5$) 79.31 0.52 0.97
-B.2 Measuring invariance of aggregate representations
-In order to have a more direct measure of invariance of the aggregate representation, we trained a model on 3DIEBench with $M_{val}=3$ and rotation conditioning, then attempted to predict the individual rotation parameters of each input view from $z_{\text{AGG}}$.
-To prevent leakage of rotation information during prediction, all action embeddings were zeroed out.
-The resulting mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; indicating that the aggregate representation contains little to no rotation-specific information.
-B.3 Complete evaluation results for linear probing on top of aggregate representations
+B.2 Results with a transformer projector instead of an MLP projector for baselines
+By using a transformer projector for baselines that normally use an MLP projector, we tested whether transformer-based projectors alone account for any performance gains.
+We did not see any benefit from switching to transformer projectors in any of the baselines.
+Table 8 shows the transformer-projector results for the baselines on 3DIEBench (plus sign indicates a transformer projector).
+Table 8: Comparison of baselines trained with transformer projectors on 3DIEBench.
+[…]
+SimCLR+ 77.92 0.32 0.48
+BYOL+ 81.05 0.09 0.16
+VICReg+ 75.10 0.15 0.28
+VICRegTraj+ 77.82 0.24 0.36
+SEN+ 81.74 0.31 0.44
+Cond. BYOL+ 73.62 0.27 0.36
+B.3 Out-of-distribution results
+We created an OOD set for 3DIEBench to enable evaluation on unseen transformations.
+The original dataset samples rotations from $(-\pi/2,\pi/2)$.
+Our OOD test set instead uses the disjoint range $(-\pi,-\pi/2)\cup(\pi/2,\pi)$, ensuring no angular overlap with the training set.
+Rendering this dataset took around five hours on a single A6000 GPU.
+From Table 9, we see that all methods fail at OOD rotation decoding (with $R^{2}$s even reaching near 1.0-1.0).
+The sharp drop in $R^{2}$ values may be attributed to a domain shift in transformation geometry: while representations are well-aligned with in-distribution rotation trajectories (inside the range $(-\pi/2,\pi/2)$), they are not geometrically structured to extrapolate beyond this range to $(-\pi,-\pi/2)\cup(\pi/2,\pi)$.
+Despite the failure in OOD equivariance generalization, seq-JEPA exhibits a graceful degradation in classification accuracy compared to other baselines.
+We hypothesize that this robustness stems from its sequence-level aggregation mechanism: even when encoder representations become less equivariant under OOD transformations, aggregation across multiple views still filters out transformation-specific variability and recovers shared semantic content.
+Thus, seq-JEPA maintains object identity better under distribution shift, indicating that aggregation over input views supports invariance-demanding downstream tasks, even when equivariance is imperfect.
+Table 9: Unseen Transformation Generalization (OOD Rotations)
+[…]
+SimCLR 63.86 (-17.27) -0.41 -0.201
+BYOL 55.68 (-27.22) -0.25 -0.198
+VICReg 61.28 (-19.20) -0.31 -0.206
+SEN 60.03 (-23.40) -0.41 -0.201
+SIE 60.19 (-17.30) -0.63 -0.200
+EquiMoD 58.54 (-25.75) -0.45 -0.206
+Cond. BYOL 57.91 (-24.70) -0.42 -0.202
+VICRegTraj 62.94 (-18.32) -0.36 -0.214
+seq-JEPA (1,1) 61.53 (-22.55) -0.69 -0.199
+seq-JEPA (3,3) 65.03 (-21.11) -0.71 -0.211
+B.4 Equivariance predictor evaluation metrics
+Following the protocol in Garrido et al.
+[2023], we compute and report MRR, Hit@1, and Hit@5 metrics to evaluate predictor quality on the 3DIEBench validation set for seq-JEPA and two predictor-based baselines (Table 10).
+These results confirm that seq-JEPA achieves strong equivariance performance not only in terms of $R^{2}$, but also in top-rank retrieval metrics.
+Table 10: Predictor evaluation metrics on 3DIEBench validation set.
+Setting MRR H@1 H@5
+SIE 0.319 0.215 0.404
+EquiMod 0.136 0.037 0.186
+seq-JEPA (1,1) 0.340 0.241 0.442
+seq-JEPA (3,3) 0.388 0.273 0.468
+B.5 Training with more compute
+To examine convergence under a high-compute regime, we trained five models using 2000 epochs, $256\times 256$ resolution, and batch size 1024.
+Each of these experiments was run on 4 A100 GPUs (our main experiments were run on a single A100).
+The evaluation results (Table 11) confirm that seq-JEPA achieves a strong performance in the high-compute regime without suffering from a trade-off between invariance and equivariance.
+Importantly, our method achieves near-saturated performance already in the low-compute regime, indicating that it requires fewer steps for convergence and is less sensitive to input resolution and batch size than competing methods.
+Table 11: Evaluation results under the high-compute regime.
+Setting Top-1 Acc $R^{2}$ (rel rot) $R^{2}$ (abs rot) MRR H@1 H@5
+SIE 82.652 0.721 0.764 0.411 0.287 0.490
+SimCLR 85.961 0.473 0.609
+EquiMod 86.833 0.492 0.625 0.154 0.048 0.201
+seq-JEPA (1,1) 85.370 0.661 0.713 0.365 0.263 0.447
+seq-JEPA (3,3) 87.581 0.736 0.781 0.419 0.282 0.483
+B.6 Complete evaluation results for linear probing on top of aggregate representations
 We provide our complete evaluation results for linear probing on top of seq-JEPA’s aggregate representations for our three transformation settings with different training and inference sequence lengths.
@@
 Figure 11: seq-JEPA’s performance on Tiny ImageNet with different types of action conditioning (crop, color jitter, blur, or all three); the heatmap shows linear probe accuracy on top of aggregate representations for different training and inference sequence lengths.
-B.4 Comparison of evaluation results on encoder representations and aggregate representations
+B.7 Comparison of evaluation results on encoder representations and aggregate representations
 For completeness, we provide linear probe classification on encoder representations for different transformation settings in Table 12 and compare them with accuracy on aggregate representations for different inference evaluation lengths.
@@
 STL-10 Position (no IoR) 4 72.49 68.95 76.84 77.97
-B.5 Additional UMAP Visualizations
+B.8 Additional UMAP Visualizations
 In Figure 12, we visualize the UMAP projections of seq-JEPA representations trained on 3DIEBench without action conditioning.
@@
 Figure 13: 2-D UMAP projections of seq-JEPA representations on STL-10 with action conditioning ($M_{tr}=M_{val}=4$).
-Encoder outputs are color-coded by class (top-left) and by $X$/$Y$ fixation coordinates (bottom); the aggregate token is color-coded by class (top-right).
+Encoder outputs are color-coded by class (top-left) and by XX/$Y$ fixation coordinates (bottom); the aggregate token is color-coded by class (top-right).
 Figure 14: 2-D UMAP projections of seq-JEPA representations on STL-10 without action conditioning ($M_{tr}=M_{val}=4$).
 Encoder outputs are color-coded by class (left) and fixation coordinates (middle); the aggregate token is color-coded by class (right).
-B.6 Details of Path Integration Experiments.
+B.9 Details of Path Integration Experiments.
 While an agent executes a sequence of actions in an environment, transitioning from an initial state to a final state, it should be capable of tracking its position by integrating its own actions.
@@
 On the other hand, ablating the visual representations (green curves) results only in a small performance drop compared to the original model, indicating that action conditioning is the key factor that enables path integration.
-B.7 Transfer learning results on ImageNet-1k
+B.10 Transfer learning results on ImageNet-1k
 To evaluate generalization of our model beyond STL-10, we assess transfer performance of the model trained on STL-10 patches via predictive learning across saccades on ImageNet-1k.
 We follow the same linear probing protocol as in-distribution evaluations: we freeze the ResNet and transformer encoder and train a linear classifier on aggregate representations generated from foveated patches.
-We extract saliency maps for ImageNet-1k images using DeepGaze IIE, resize them to $224\times 224$, and sample sequences of patches sized $32\times 32$ or $84\times 84$ for this out-of-distribution (OOD) evaluation setting.
+We extract saliency maps for ImageNet-1k images.
+Saliencies were extracted using DeepGaze IIE with the MIT1003 centerbias prior [Linardos et al., 2021].
+Maps are saved at native resolution (matching the original ImageNet image dimensions) and normalized to probability distributions.
+For our transfer learning experiments we resize the images and saliencies to $224\times 224$, and sample sequences of patches sized $32\times 32$ or $84\times 84$ for this evaluation setting.
 Figure 15 shows top-1 linear probe accuracy on ImageNet-1k validation set across varying inference sequence lengths ($M_{val}$).